### v1

# Neuromodulation and homeostasis: complementary mechanisms for robust neural function

ABSTRACT
Neurons depend on two interdependent mechanisms—homeostasis and neuromodulation—to maintain robust and adaptable functionality. Homeostasis stabilizes neuronal activity by adjusting ionic conductances, whereas neuromodulation dynamically modifies ionic properties in response to external signals. Combining these mechanisms in conductance-based models often produces unreliable outcomes, particularly when sharp neuromodulation interferes with homeostatic tuning. This study explores how a biologically inspired neuromodulation controller can harmonize with homeostasis to ensure reliable neuronal function. Using computational models of stomatogastric ganglion and dopaminergic neurons, we demonstrate that controlled neuromodulation preserves neuronal firing patterns while maintaining intracellular calcium levels. Unlike sharp neuromodulation, the neuromodulation controller integrates activity-dependent feedback through mechanisms mimicking G-protein-coupled receptor cascades. The interaction between these controllers critically depends on the existence of an intersection in conductance space, representing a balance between target calcium levels and neuromodulated firing patterns. Maximizing neuronal degeneracy enhances the likelihood of such intersections, enabling robust modulation and compensation for channel blockades. We further show that this controller pairing extends to network-level activity, reliably modulating central pattern generators in crustaceans. These findings suggest that targeting neuromodulation pathways—rather than ion channels directly—may offer safer pharmacological strategies to manage neuronal dysfunctions. This study highlights the complementary roles of homeostasis and neuromodulation, proposing a unified control framework for maintaining robust and adaptive neural activity under physiological and pathological conditions.
SIGNIFICANCE STATEMENT
Neurons must maintain stable activity while adapting to changing demands, relying on two mechanisms: homeostasis, which stabilizes activity, and neuromodulation, which adjusts behavior to external signals. These mechanisms often interact, but their improper coordination can lead to dysfunction. This study uses computational models to show how controlled neuromodulation - mimicking biological feedback - can be harmonized with homeostasis to ensure robust neuronal function. By improving neuronal degeneracy, the system compensates for channel blockades and maintains critical activity patterns for both single neurons and networks. These insights not only deepen our understanding of neural robustness, but also suggest safer pharmacological strategies targeting neuromodulatory pathways, offering potential breakthroughs in treating neurological disorders without disrupting essential cellular mechanisms.
INTRODUCTION
Brain activity is continuously shaped by neuromodulators and neuropeptides, including dopamine, serotonin, and histamine (Bargmann and Marder,, 2013; Marder et al.,, 2014). Neuromodulators dynamically influence single-neuron activity, input-output properties, and synaptic connection strength and dynamics, enabling neuronal networks to adapt to changing needs, contexts, and environments (Marder and Calabrese,, 1996; Marder and Bucher,, 2001; McCormick et al.,, 2020). To achieve this modulation, neuromodulators dynamically reshape the density, dynamics, and kinetics of many single cell transmembrane proteins, making whole brain functional signaling strongly dependent on the robustness and reliability of neuromodulation mechanisms at the molecular and cellular levels (Nadim and Bucher,, 2014; Hooper and Marder,, 1984). While experimental studies underscore the ubiquity of neuromodulation across nervous systems, the fundamental principles underlying this mechanism remain incompletely understood. As such, studying neuromodulation is a very arduous task, both experimentally and computationally, which explains our limited understanding of its underlying mechanisms.
Neuromodulation coexists with cellular homeostasis, a mechanism that gradually adjusts neuronal membrane properties to maintain a target activity level (Desai,, 2003; Marder and Goaillard,, 2006; O’Leary et al.,, 2013). Neuronal cells are dependent on homeostatic mechanisms to maintain optimal functionality throughout the extended useful life of mammals (Marder,, 2011). Despite continuous turnover of transmembrane proteins, such as ion channels and receptors, occurring on varying temporal scales ranging from hours to weeks (O’Leary and Wyllie,, 2011), the robustness and excitability of neurons must remain undisturbed (Marder and Prinz,, 2002). By monitoring only intracellular calcium levels, the neuron orchestrates the regulation of all its ion channel conductances through homeostatic self-tuning rules (O’Leary et al.,, 2010; Pratt and Aizenman,, 2007). This raises the question of how two different mechanisms acting on the same targets (ion channels), each potentially having divergent goals, can co-exist and complement each other harmoniously (Marder et al.,, 2014; Marder and Prinz,, 2002).
We propose to study this interaction between neuromodulation and homeostasis computationally, using conductance-based models. Conductance-based models provide an electrical representation of the neuronal membrane, where the membrane is modeled as an RC circuit, while the ion channels are modeled as non-linear conductances (Izhikevich,, 2007). These models are the closest equivalent of biological neurons, thanks to their ability to model any ion channel from experimental data following a methodology first proposed by Hodgkin and Huxley (Hodgkin and Huxley,, 1952). There exists a myriad of different conductance-based models that model various types of neuronal cells (Liu et al.,, 1998; Qian et al.,, 2014; Plant,, 1981; Wechselberger et al.,, 2006; Drion et al.,, 2011; Katori et al.,, 2010; Megwa et al.,, 2023).
On the one hand, a homeostatic controller has been developed by O’Leary et al., (2014) and can be applied on top of any conductance-based model. In this controller, the neuron can adjust all ionic conductances up or down depending on its average intracellular calcium level. If the calcium level is lower than a reference value, ionic conductances are increased following homeostatic tuning rules and vice versa. Such rules force the tuning of ionic conductances along a single axis, that is, by preserving conductance ratios or correlations. On the other hand, Fyon et al., (2023) has developed a reliable neuromodulation controller and can also be applied to any conductance-based model. This controller can robustly tune the external activity of neurons, which means that neurons with various sets of ion channel properties that lead to similar firing activities (a property called degeneracy (Drion et al., 2015b, )) can be neuromodulated using the same control action. In this study, we first examine the interaction between homeostasis and sharp neuromodulation, emphasizing that improperly combining the homeostatic controller with neuromodulation may result in unreliable outcomes. Subsequently, we expand our analysis to explore the interaction between the homeostatic controller and a neuromodulation controller, both at the single-cell level and in rhythmic networks. This approach enables the development of a combined controller that integrates the strengths of both individual controllers, facilitating robust modulation and preservation of neuronal function.
RESULTS
Current model of homeostasis can lead to pathological excitability in the presence of sharp neuromodulation
This study employs a stomatogastric ganglion (STG) neuron conductance model (Liu et al.,, 1998) to explore the interaction between homeostasis and neuromodulation. These neurons contribute to rhythmic circuits that generate signals triggering stomach muscle contractions in crustaceans. This model was chosen for its ability to produce diverse firing patterns—ranging from tonic spiking to bursting, including single-spike bursting—and its rich combination of ion channels enabling degeneracy. The model includes classical fast sodium and delayed rectified potassium channels, A-type and calcium-activated potassium channels, T-type and slow calcium channels as well as hyperpolarization-activated H-channels. We first study the interaction between state-of-the-art model homeostasis (O’Leary et al.,, 2014) and sharp neuromodulation (Fyon et al.,, 2024). We define sharp neuromodulation as a direct change in a subset of ion channel maximal conductances, which results in functional modulation. In the context of this work, we will focus on a robust transition from tonic spiking to bursting in degenerate neurons, i.e. neurons exhibiting similar firing activities from different underlying channel conductances.
First, we sharply neuromodulated a set of degenerate neurons from tonic spiking to bursting by directly modifying the maximal conductances of A-type potassium $\bar{g}_{A}$ and slow calcium $\bar{g}_{CaS}$, then allowed homeostatic compensation to take its course. Three example traces with a time evolution of maximal conductances are provided in Fig. 1, left, whereas Fig. 1, right show trajectories in the $\bar{g}_{\mathrm{CaS}}$-$\bar{g}_{\mathrm{A}}$ plane. As can be seen in the figure, although the initial tonic firing and the sharply neuromodulated bursting activities are very similar in all three neurons, homeostatic compensation has very variable and unreliable effects. Depending on the initial conductance state, neuromodulation shifts the neuron to a new state where homeostatic compensation can lead to pathological (Fig. 1A), functionally preserved (Fig. 1B), or unphysiological (Fig. 1C) behaviors. This interesting result highlights an important property of neuronal degeneracy: although distinct channel subsets can produce similar neuronal behaviors, neurons may respond differently to perturbations, such as sharp neuromodulation.
To better understand the reasons behind this unreliability, both mechanisms must be discussed in detail. In the homeostatic model O’Leary et al., (2014), the cell senses its own intracellular calcium level and averages it over time. If the mean intracellular calcium falls below a predefined target value, RNA transcription and ion channel protein translation are triggered to increase all ionic conductance values, and vice versa. Specifically, increasing calcium ion channel conductances leads to more calcium ions entering the cell. This implies that homeostasis induces changes in all conductance values (Fig. 1 left), which leads to a linear motion in the conductance plane (Fig. 1 right). This phenomenon happens on a relatively slow timescale, which can last up to days (O’Leary et al.,, 2010). Notably, the RNA transcription time constants, denoted by $\tau_{m}$, vary across different ion channels and cell types. Steady-state mathematical analysis has shown that these differing time constants tune correlations in ion channel conductances. For two ion channels $i$ and $j$, steady-state analysis yields $\bar{g}_{i}/\bar{g}_{j}\approx\tau_{m_{j}}/\tau_{m_{i}}$. This indicates that homeostatic tuning rules maintain ratios of ionic conductances and their correlations while adjusting their values in response to intracellular calcium levels. In the ionic conductance space, these rules allow neurons to navigate along a one-dimensional line, where the slope of this line in two dimensions is determined by the ratios of mRNA transcription time constants. This phenomenon, known as homogeneous scaling, involves proportionally increasing or decreasing all ionic conductances by the same multiplicative factor.
On the other hand, sharp neuromodulation is implemented by transiently applying the algorithm of Fyon et al., (2024). This neuromodulation also produces changes in conductance values (Fig. 1 left), resulting in motion within the conductance plane in a direction different from homogeneous scaling (Fig. 1 right) to reach a new firing activity. However, neuromodulation only affects a subset of ion channels, not all (Marder,, 2012). Furthermore, the timescale of neuromodulation is much faster than that of homeostasis, lasting for a maximum of a few minutes (Marder et al.,, 2014). In addition, neuromodulation tunes correlations between neuromodulated ion channels, rather than preserving them (Fyon et al.,, 2024).
Coupling homeostasis and neuromodulation in this way is problematic because homeostasis preserves correlations and regulates calcium, whereas neuromodulation adjusts correlations to modulate neuronal function. In Fig. 1, sharp neuromodulation is applied to three different STG models with distinct initial ionic conductances. Initially, all neurons exhibited tonic firing due to homeostatic regulation, with RNA transcription time constants selected to produce tonic spiking neurons. Neuromodulation was then applied instantaneously, leading to a transition from tonic spiking to bursting in all cases (Fig. 1). These transitions drastically increase intracellular calcium concentration. From the perspective of homeostatic tuning rules, this constitutes a disturbance, as it alters the conductance ratios between $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$. Although this disturbance is necessary for adaptation and functional modulation, the outcome of homeostasis in response to this biological disturbance is unpredictable. Indeed, as the homeostatic controller was initially designed to produce tonic spiking neurons, it tends to preserve the predefined correlations that produce such behavior. However, neuromodulation changes the firing activity, leading to a mismatch between actual conductance values and the encrypted conductance ratios.
In Fig. 1A, homeostatic compensation results in pathological behavior, characterized by irregular bursting with a hyperpolarized plateau, failing to achieve either bursting or tonic spiking. In Fig. 1B, the modulated behavior (bursting) is maintained after homeostatic compensation. In Fig. 1C, homeostatic compensation leads to unphysiological values of $\bar{g}_{\mathrm{A}}$, tending towards negative values. Even when $\bar{g}_{\mathrm{A}}$ is constrained to be non-negative, the resulting electrical behavior is pathological, failing to exhibit either tonic spiking or bursting. This toy experiment highlights that homeostasis and sharp neuromodulation can interfere pathologically.
Such behavior was predicted in O’Leary et al., (2014), where the authors demonstrate that their homeostasis model can lead to various outcomes in response to disturbances, such as channel deletions. Initially, the deletion itself may directly affect neuronal function, either preserving it or causing its loss. Subsequently, homeostatic compensation can either preserve or restore function, but it may also lead to pathological behaviors, such as functional deletion. This result suggests that any disturbance applied to the conductance ratios—even physiological ones such as neuromodulation—can lead to problematic behaviors.
Figure 1: Current model of homeostasis can lead to pathological excitability in the presence of sharp neuromodulation. A. Time evolution of all conductances in the STG model, shown on a logarithmic scale, when homeostasis is applied with sharp neuromodulation (instantaneous changes in $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$ within the red region) (left). Corresponding trajectory in the modulated conductance space (right). The white circle represents the initial condition (tonic spiking), the red diamond marks the neuromodulated state (bursting) immediately after the sharp modulation, and the green triangle denotes the final state after homeostasis has converged. In this case, homeostatic compensation is pathological, resulting in irregular bursting. B. Same as in panel A, but with different degenerate initial conditions. Here, homeostatic compensation preserves the neuromodulated function. C. Same as in panel A, with another set of degenerate initial conditions. In this case, homeostatic compensation results in negative values of $\bar{g}_{\mathrm{A}}$ (saturated at 0), leading to pathological behavior characterized by one-spike bursting.
Controlled neuromodulation and homeostasis naturally lead to robust neuronal function
In biological systems, homeostasis and neuromodulation collaborate to ensure robust and modular neural function. These mechanisms operate on different timescales: neuromodulation acts within seconds or minutes (Werner and Mitterauer,, 2014), while homeostatic compensation may take hours or even days (Marder and Prinz,, 2002). However, in computational models, the goals of neuromodulation and homeostatic controllers often seem discordant. Neuromodulation typically adjusts conductance ratios, while homeostasis seeks to maintain these ratios and guide the neuron through homogeneous scaling. Consequently, applying sharp neuromodulation to any homeostatic conductance-based model can lead to instability, as neuromodulation disrupts the references and control actions of the homeostasis controller (Fig. 2A).
To model the actions of both mechanisms within a single cell, we include a more advanced neuromodulation model, referred to as the neuromodulation controller, as developed in Fyon et al., (2023). This controller can be coupled with any conductance-based model, allowing modulation of its neuronal function, such as its firing pattern, to mimic biological neuromodulation. We call this type of neuromodulation controlled neuromodulation, in contrast to the sharp neuromodulation discussed above. The neuromodulation controller adjusts target channel conductances, providing continuous and consistent modulation rather than instantaneous, rigid changes.
Fig. 2B illustrates the results of transitions from tonic spiking to bursting when applied to a degenerate population of STG neuron models. Initially, all neurons exhibit tonic firing mode, and homeostasis adjusts the conductances to achieve a predefined calcium level along a first tonic spiking homogeneous scaling line. Midway through the simulation, controlled neuromodulation is applied to switch all neurons to bursting mode, which now occurs through a controlled adjustment of $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$. Consequently, intracellular calcium levels rise sharply, and conductance ratios involving one or both neuromodulated ionic conductances are adjusted. Finally, with the reshaped RNA time constants, the homeostasis controller adjusts all conductances to restore an appropriate calcium level, while controlled neuromodulation preserves neuromodulated conductance ratios along a second bursting homogeneous scaling line. Ultimately, the neuronal population reaches a new state where the neuromodulated function is preserved and intracellular calcium levels are back to their target value.
Simulations were also performed on a midbrain dopaminergic neuron model (adapted from Qian et al., (2014)) to demonstrate the generality of our results (Fig. 2C). As with the previous simulations, a transition from tonic spiking (pacemaking) to strong bursting was induced midway through by modulating the N-type and L-type calcium conductances ($\bar{g}_{\mathrm{CaN}}$ and $\bar{g}_{\mathrm{CaL}}$) using controlled neuromodulation. Homeostatic compensation was continuously applied to maintain proper calcium levels, resulting in a reliable and robust outcome despite initial degeneracy within the population.
Figure 2: Controlled neuromodulation and homeostasis naturally lead to robust neuronal function. A. Schematic representation of neuromodulation (left) and homeostasis (right) cascades. Neuromodulators bind to G-protein receptors, triggering complex signaling pathways that selectively affect subsets of ion channels. In contrast, homeostasis senses intracellular calcium levels and regulates them by adjusting all conductances uniformly. B. Time evolution of all conductances in the STG model, displayed on a logarithmic scale, during homeostasis with controlled neuromodulation (controlled changes in $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$ starting at the dashed blue line) for a degenerate population of 200 models (left). The corresponding mean intracellular calcium concentration over time (right) shows the target value (red dash-dotted line) is maintained both before and after controlled neuromodulation. C. Same as panel B, but for a DA model where controlled neuromodulation affects $\bar{g}_{\mathrm{CaN}}$ and $\bar{g}_{\mathrm{CaL}}$.
Why do controlled neuromodulation and homeostasis lead to robust neural function?
Fig. 3A shows the trajectories of the simulations from Fig. 2B in the plane of modulated conductances, $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$, during tonic spiking homeostatic compensation for both sharp and controlled neuromodulation. Initially, the modulated conductances of every neuron in the population grow along the homogeneous scaling direction in tonic spiking mode, as the calcium levels are below the threshold value. This homogeneous scaling corresponds to a line intersecting the origin, with variability due to the degeneracy in conductance ratios within the population. During neuromodulation, all conductances are adjusted to tune their ratios and achieve bursting (Fig. 3B). Importantly, all neuromodulatory actions occur along the same direction, but with varying magnitudes depending on the state of the neuron (Fyon et al.,, 2024). As a result, calcium levels increase drastically for each neuron in the population. Once the neuromodulatory action ends, the outcome depends on the type of neuromodulation applied (Fig. 3C).
In the case of sharp neuromodulation, during the homeostatic action in the bursting phase, the homeostatic controller acts alone and tends to maintain the same conductance ratios as in tonic spiking (Fig. 3C left). This leads to a reduction in calcium levels, but, as seen in Fig. 1, its effects are unreliable and undesirable. Specifically, for most neurons in the population, the homeostatic controller reduces $\bar{g}_{\mathrm{A}}$ towards negative values (saturated at 0 to avoid model instability) to bring $\bar{g}_{\mathrm{CaS}}$ into an appropriate range. This regulation causes the neuron to move away from the robust bursting behavior, resulting in heterogeneous changes in the activity of the population.
In the case of controlled neuromodulation, the homeostatic and neuromodulation controllers work together to maintain the new conductance ratios corresponding to bursting (Fig. 3C right). This results in a new homogeneous scaling following these updated conductance ratios in bursting, which translates into trajectories that move toward the origin of the axes, as would be expected in tonic spiking. Homogeneous scaling during bursting differs across the population due to initial degeneracy. The difference between the two homogeneous scalings (tonic spiking and bursting) is driven by the neuromodulatory action, ensuring that the neuromodulated behavior (bursting) is not lost.
Figure 3: Why do controlled neuromodulation and homeostasis lead to robust neural function? A. Trajectories of the STG population from Fig. 2 in the modulated conductance space during homogeneous scaling in tonic spiking only. As expected from homeostasis, all neurons move in the same direction. White circles represent the initial conditions at the start of the step, and the green triangle marks the final conditions. B. Same as panel A, but during the brief period when neuromodulation is active. All trajectories are parallel to one another. C. Same as panels A and B, but during the homeostasis phase following neuromodulation. The outcome depends on the type of neuromodulation: sharp (left) or controlled (right). Sharp neuromodulation results in non-robust outcomes, whereas controlled neuromodulation preserves robust neural function.
In the first section, when the homeostasis controller was combined with sharp neuromodulation, unreliable outcomes were observed. To explain this behavior, we use a schematic representation of the modulated conductance plane (Fig. 4A). In this plane, different values of $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$ can lead to similar mean intracellular calcium concentration. These values can be highlighted using calcium isoclines (green dashed lines in Fig. 4A). Likewise, different values of $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$ can lead to similar firing activity. These values can be highlighted using activity isoclines (blue dashed lines in Fig. 4A). Calcium isoclines are parallel one to each other, whereas activity isoclines pass through the origin and rotate around this point. This behavior is explained by homeostatic tuning rules and homogeneous scaling, as different conductance ratios lead to different firing patterns. Initially, the neuron operates in the target spiking mode at the target calcium level, corresponding to the intersection of the spiking and the target calcium isoclines. When sharply neuromodulating the neuron, the modulated conductances change, and the neuron shifts to another state, now reaching a new target strong bursting isocline, which intersects a higher calcium isocline. Homeostasis then kicks in to reduce the calcium level to its target value. However, in the homeostatic controller, the “native” tonic spiking direction is encrypted, leading to changes in conductances along the spiking isocline. As a result, the neuron moves away from the bursting isocline, potentially disrupting the new function. This explains why sharp neuromodulation is incompatible with the homeostasis controller.
To harmonize the actions of both controllers and achieve robust functional modulation, we leverage the significant difference in their timescales. Neuromodulation controls the modulated ionic conductances, while homeostatic tuning rules regulate all ionic conductances over a slower timescale to maintain an appropriate calcium level. In the modulated conductance plane (Fig. 4B), the initial and neuromodulated states are the same as in Fig. 4A. However, in the case of controlled neuromodulation, both homeostatic and neuromodulatory controllers are active. Specifically, when homeostasis reduces the calcium level and shifts the neuron away from the bursting isocline, the neuromodulation action counteracts this perturbation in function by bringing the neuron back to the bursting isocline, preserving its function. The neuromodulatory controller operates several orders of magnitude faster than the homeostatic controller. As a result, fast corrections from the former render perturbations from the latter infinitesimal, leading to changes in the conductance ratio along the target activity isocline (strong bursting in this example). Ultimately, the neuron stabilizes at the intersection of the target calcium and target activity isoclines, where both controllers meet their objective.
Figure 4: Schematic explanation of why controlled neuromodulation integrates better with homeostasis compared to sharp neuromodulation. A. Schematic of the modulated conductance space for homeostasis combined with sharp neuromodulation. Initially (white circles), the neuron spikes at the target calcium level, located at the intersection of the firing pattern isocline (dashed blue lines) and the calcium level isocline (dashed green lines). Following sharp neuromodulation to a new firing pattern (red diamond), calcium levels sharply increase. Homeostasis reduces calcium (solid red arrow) by moving along the same direction as before neuromodulation (dashed red arrow), causing the neuron to deviate from the neuromodulated firing pattern isocline. B. Same as panel A, but with controlled neuromodulation. The initial steps are identical, but after neuromodulation, as homeostasis acts to reduce calcium, controlled neuromodulation simultaneously adjusts to keep the neuron on the neuromodulated firing pattern isocline (sawtooth red arrow). Because controlled neuromodulation operates on a much faster timescale than homeostasis, the neuron remains on the neuromodulated firing pattern isocline throughout.
Controlled neuromodulation and homeostasis ensure the preservation of function under physiologically recoverable disturbances
By employing this tandem of controllers, neuronal function becomes robust to channel blockade, provided that the channel deletion is compensable. While homeostasis compensation alone might lead to unreliable recovery from channel blockade, incorporating controlled neuromodulation somehow stabilizes the neuron response to such blockade, even in the case of highly degenerate neuronal populations. Furthermore, if the channel blockade is physiologically recoverable—meaning the function of the blocked channel can be restored using other channels—there exists an optimal combination of neuronal feedback gains to preserve that function. In essence, if one channel is blocked, a specific concentration of particular neuromodulators can preserve the function.
For example, in Fig.5, various channel blockades are applied to a degenerate population of STG conductance-based models equipped with the combined homeostasis/neuromodulation tandem of controllers. The target behavior of the population is regular bursting. In Fig.5A, H-type channels are blocked. Here, the function is preserved even immediately after the blockade, without requiring compensation. Subsequently, the combined regulation of homeostasis and controlled neuromodulation maintains this function by restoring the neuron to an appropriate calcium level on a population-wide scale. In Fig. 5B and C, T-type calcium and calcium-controlled potassium channels are blocked, respectively. In these cases, the blockade results mostly in the loss of bursting function. However, by allowing the homeostasis/neuromodulation pair of controllers to compensate, the function is largely recovered, albeit with irregular bursting patterns in some neurons. Additionally, intracellular calcium levels stabilize at their target values. This preservation of functions is feasible because the actions of H-type, T-type calcium, and calcium-controlled potassium channels are degenerate and can be compensated for using other channels present within the cell. The STG neuron model incorporates numerous channels operating on slow and ultraslow timescales, enabling the compensation for lost channels in these timescales. It is worth noting that only the loss of the calcium-controlled potassium channel has been compensated using different neuronal feedback gain values, i.e., different concentrations of various neuromodulators, as in control conditions.
However, the blockade of fast transient sodium channels is irrecoverable (Fig. 5D), as these channels are crucial for initiating spikes on a fast timescale. Once the blockade is applied, compensation cannot restore spikes because the action of sodium channels is non-degenerate and occurs on the fast timescale of the neuron. Similarly, achieving the appropriate intracellular calcium level becomes unattainable. A comparable outcome would result from the blockade of delayed rectified potassium channels, which are responsible for the downstroke of the spikes.
Figure 5: Controlled neuromodulation and homeostasis ensure the preservation of function under physiologically recoverable disturbances. A. Time evolution of all conductances in the STG model, displayed on a logarithmic scale, during homeostasis with controlled neuromodulation (controlled changes in $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$ starting at the dashed blue line) for a degenerate population of 200 models with H-type ion channel blockade (left). The corresponding distribution of behaviors before the blockade, immediately after the blockade, and following compensation (right) demonstrates the effect of the blockade. In this case, no neurons lose bursting due to the channel blockade or compensation. B. Same as panel A, but with a CaT channel blockade. The blockade causes half of the neurons to lose bursting; however, all neurons recover bursting with compensation. C. Same as panel A, but with a KCa channel blockade. The blockade causes more than half of the neurons to lose bursting, but nearly all recover bursting with compensation. D. Same as panel A, but with a Na channel blockade. The blockade causes all neurons to lose bursting, and compensation does not restore bursting due to the non-degenerate essential role of Na channels.
Central pattern generator networks can be robustly modulated to orchestrate their rhythmic output
The application of this tandem of controllers extends even to networks, particularly central pattern generators (CPGs). CPGs are self-organized neuronal circuits mainly composed of bursting neurons connected primarily through inhibitory connections, capable of generating rhythmic output without external input (Yuste et al.,, 2005; Guertin,, 2009). One extensively studied circuit is responsible for the pyloric and gastric mill rhythms in crustaceans, observed in the stomatogastric ganglion. This network exhibits two rhythms at different speeds: the fast pyloric rhythm and the slow gastric mill rhythm. These rhythmic outputs are vital for coordinating muscle contractions in the stomach for digestion, among other functions. A minimalist network model of such a bi-rhythmic circuit is investigated in Fig. 6 (Gutierrez et al.,, 2013; Drion et al.,, 2019). It comprises two half-center oscillators operating at different speeds, connected via a central neuron. Half-center oscillators consist of two bursting neurons connected in mutual inhibition, enabling them to fire anti-phase bursts. In the network scheme, the purple and blue neurons constitute the fast half-center oscillator (representing the pyloric rhythm), while the green and yellow neurons form the slow half-center oscillator (representing the gastric mill rhythm). The red neuron serves as the central one, also in bursting mode. It has been demonstrated that the gastric mill rhythm requires a neuromodulatory input to be activated (Selverston et al.,, 2009).
In this study, the gastric mill rhythm can be reliably induced by the pair of controllers, but not with homeostasis combined with sharp neuromodulation. Under control conditions (Fig. 6A), i.e., without neuromodulatory input to the gastric mill rhythm, only the pyloric rhythm is active and the central neuron consistently bursts. In particular, variability in conductances is observed in this network (Fig. 6A right). When neurons are equipped with homeostasis only, activation of sharp neuromodulation for the gastric mill rhythm leads to rapid loss of function, due to the unreliability of this combination (Fig. 6B middle). However, when neurons undergo controlled neuromodulation, the gastric mill rhythm is successfully induced and sustained (Fig. 6B right). Despite the degeneracy of neurons, characterized by significant variability in conductance ratios, the network achieves rhythmic output with appropriate intracellular calcium levels. This toy experiment demonstrates that such a tandem of controllers can be applied in computational models to modulate and maintain not only single cell activity but also network-level activity.
Figure 6: Central pattern generator networks can be robustly modulated to orchestrate their rhythmic output. A. Control condition of the simplified pyloric/gastric mill rhythm network (left), inspired from Gutierrez et al., (2013). Without neuromodulation, the fast rhythm is active (top two traces, middle), while the slow rhythm is inactive (bottom two traces, middle). In this case, all neurons in the network exhibit degenerate conductances (right). B. Schematic representation of the activated slow rhythm network (left), modulated by either sharp neuromodulation (middle) or controlled neuromodulation (right). With sharp neuromodulation, the slow rhythm fails due to its unreliability, whereas controlled neuromodulation sustains the slow rhythm.
DISCUSSION
The efficient and robust integration of homeostasis and neuromodulation in neurons remains an open question with several proposed hypotheses (Marder et al.,, 2014; Marder and Bucher,, 2005; Cunha,, 2001). The improper integration of these mechanisms—for example, combining sharp neuromodulation with a homeostatic controller—can lead to negative interactions. Such interactions may result in unreliable behaviors or numerical instabilities.
In contrast, pairing a homeostatic controller with controlled neuromodulation ensures reliable neural function modulation while maintaining appropriate calcium levels. This approach draws inspiration from biology, where neuromodulation is known to be activity-dependent, as observed experimentally (Kramer and Levitan,, 1990; Raymond et al.,, 1992). Specifically, the neuromodulation controller models the role of G-protein-coupled receptors and second messengers, embedding feedback in the metabolic neuromodulation cascade (Marcus and Bruchas,, 2023; Scheler,, 2004). This reliable interaction can be extended to neural circuits, such as the gastric mill circuit in crabs, to produce robust network-level functionality (Marder et al.,, 2014).
However, the success of this interaction critically depends on a point in the conductance space where both controllers converge. This point represents a neural activity pattern that supports neuromodulated firing while maintaining homeostatic calcium levels. As shown in Fig. 4, this intersection of the strong bursting isocline (the target of controlled neuromodulation) and the target calcium level (the goal of homeostatic control) represents the convergence point. When such a point is absent—e.g., during sodium channel blockade (Fig. 5)—neither neuromodulated activity nor calcium homeostasis is sustained. This may lead to pathological compensations, where one controller prioritizes calcium regulation at the cost of abandoning the desired firing pattern (Starmer et al.,, 1991; Grant et al.,, 1984; Anger et al.,, 2001). Furthermore, discontinuities in the neuromodulated firing pattern may lead to transient pathological behaviors.
Maximizing convergence between these control mechanisms requires increasing the likelihood of both an intersection and a continuous path in conductance space. We hypothesize that this can be achieved by maximizing neuronal degeneracy (Whitacre,, 2010; Rathour and Narayanan,, 2019; Whitacre and Bender,, 2010; Prinz,, 2017; Goaillard and Marder,, 2021). Higher degeneracy increases the likelihood that the functional and dynamical role of one ion channel can be compensated by others, thereby preserving both the path and the intersection. This underscores the essential role of degeneracy in enabling robust neural functions that can be modulated.
Finally, this study underscores the potential risks associated with pharmacological ion channel blockade (X Cubeddu,, 2016; DeWitt and Waksman,, 2004; Imbrici et al.,, 2018). If neuronal degeneracy is insufficient, such interventions may disrupt neuron ability to maintain robust neuromodulation. Instead, we propose targeting elements within the neuromodulation cascade—such as second messengers or other components of the control loop—as more reliable pharmacological strategies. Such interventions would preserve the dynamics of controlled neuromodulation, providing a more robust alternative to direct ion channel blockade.
MATERIALS AND METHODS
Programming language
The Julia programming language was used in this work Bezanson et al., (2017). Numerical integration was realized using DifferentialEquations.jl.
Conductance-based models
For all experiments, single-compartment conductance-based models were employed. These models articulate an ordinary differential equation for the membrane voltage $V$, where $N$ ion channels are characterized as nonlinear dynamic conductances, and the phospholipid bilayer is represented as a passive resistor-capacitance circuit. Mathematically, the voltage-current relationship of any conductance-based neuron model is expressed as follows:
$I_{C}$ $=C\frac{\mathrm{d}V}{\mathrm{d}t}+g_{\mathrm{leak}}(V-E_{\mathrm{leak}})=-I_{\mathrm{int}}+I_{\mathrm{ext}},$
$=-\sum_{\mathrm{ion}\in\mathcal{I}}g_{\mathrm{ion}}(V,t)(V-E_{\mathrm{ion}})+I_{\mathrm{ext}}.$
Here, $C$ represents the membrane capacitance, $g_{\mathrm{ion}}$ denotes the considered ion channel conductance and is non-negative, gated between 0 (all channels closed) and $\bar{g}_{\mathrm{ion}}$ (all channels open), $E_{\mathrm{ion}}$ and $E_{\mathrm{leak}}$ are the channel reversal potentials, $\mathcal{I}$ is the index set of intrinsic ionic currents considered in the model, and $I_{\mathrm{ext}}$ is the current externally applied in vitro, or the combination of synaptic currents. Each ion channel conductance is nonlinear and dynamic, represented by $g_{\mathrm{ion}}(V,t)=\bar{g}_{\mathrm{ion}}m_{\mathrm{ion}}^{a}(V,t)h_{\mathrm{ion}}^{b}(V,t)$, where $m_{\mathrm{ion}}$ and $h_{\mathrm{ion}}$ are variables gated between 0 and 1, modeling the opening and closing gates of ion channels, respectively. Throughout this study, both the isolated crab STG neuron model Liu et al., (1998) and the adapted DA neuron model Qian et al., (2014) (where SK channels had been blocked to enable bursting) were employed.
The STG model consists of seven ion channels that operate on various time scales: fast sodium channels ($\bar{g}_{\mathrm{Na}}$); delayed-rectifier potassium channels ($\bar{g}_{\mathrm{Kd}}$); T-type calcium channels ($\bar{g}_{\mathrm{CaT}}$); A-type potassium channels ($\bar{g}_{\mathrm{A}}$); slow calcium channels ($\bar{g}_{\mathrm{CaS}}$); calcium controlled potassium channels ($\bar{g}_{\mathrm{KCa}}$); and H channels ($\bar{g}_{\mathrm{H}}$).
The DA model consists of six ion channels that operate on various time scales: fast sodium channels ($\bar{g}_{\mathrm{Na}}$); delayed-rectifier potassium channels ($\bar{g}_{\mathrm{Kd}}$); L-type calcium channels ($\bar{g}_{\mathrm{CaL}}$); N-type calcium channels ($\bar{g}_{\mathrm{CaN}}$); ERG channels ($\bar{g}_{\mathrm{ERG}}$); and NMDA channels ($\bar{g}_{\mathrm{NMDA}}$). Note that, owing to the multicellular nature of NMDA channels, they were excluded from this study but were still used for simulations with baseline values.
The homeostatic controller
The homeostatic controller adjusts all conductances using an integration rule described in O’Leary et al., (2014). Let $\bar{g}_{i}$ denote the conductance of ion channel $i\in[1,N]$, and $m_{i}$ the corresponding mRNA concentration. The dynamics of the homeostatic controller are governed by the following equations:
$\tau_{i}\dot{m}_{i}$ $=[Ca^{+2}]_{\mathrm{target}}-[Ca^{+2}],$
$\tau_{g}\dot{\bar{g}}_{i}$ $=m_{i}-\bar{g}_{i},$
where $[Ca^{+2}]$ and $[Ca^{+2}]_{\mathrm{target}}$ represent the intracellular calcium concentration and its target value, respectively. In essence, the calcium concentration error is integrated into the mRNA levels, which subsequently modulate the corresponding channel conductances. This regulation aims to bring the calcium concentration to its target value, as higher channel conductances (particularly calcium-conducting ones) facilitate greater calcium influx. Starting from any initial condition, this controller drives the cell along a line in conductance space known as the homogeneous scaling line. This line passes through both the initial condition and the origin of the conductance space axes. Steady-state analysis reveals that the conductance ratios remain invariant, such that $\frac{\bar{g}_{i}}{\bar{g}_{j}}=\frac{\tau_{j}}{\tau_{i}}$. This invariance ensures that the relative scaling of conductances is preserved during homeostatic compensation. Consequently, the cell’s conductance properties develop along the homogeneous scaling beam defined by these ratios.
The neuromodulation controller
The neuromodulation controller adjusts a subset of $n$ conductances, denoted as $\bar{g}_{\mathrm{mod}}\in\mathbb{R}^{n}$ with $n<N$, using the algorithm introduced in Fyon et al., (2023). This controller employs the concept of Dynamic Input Conductances (DICs) to determine target values for $\bar{g}_{\mathrm{mod}}$ based on the input neuromodulator concentration, represented as $\bar{g}_{0}\left([\mathrm{nmod}]\right)\in\mathbb{R}^{n}$. The error between the target and the current values, $e_{\mathrm{mod}}$, drives a Proportional-Integral (PI) controller, which updates $\bar{g}_{\mathrm{mod}}$ to track the desired reference. The controller dynamics are described by the following equations:
$e_{\mathrm{mod}}$ $=\bar{g}_{0}\left([\text{nmod}]\right)-\bar{g}_{\mathrm{mod}},$
$\dot{\bar{g}}_{\mathrm{mod}}$ $=f\left(K_{p}\cdot e_{\mathrm{mod}}+K_{i}\cdot\int e_{\mathrm{mod}}\mathrm{d}t\right),$
where $K_{p}$ and $K_{i}$ are the proportional and integral gains of the PI controller, respectively. In summary, the controller adjusts the modulated conductances $\bar{g}_{\mathrm{mod}}$ in response to the local concentration of neuromodulator, producing new reference values for these conductances to fine-tune the cell’s firing pattern—effectively neuromodulating it. This neuromodulator controller produces changes in conductance ratios that are normally preserved within the homeostatic controller. The new reference values depend on the neuromodulator concentration, the type of neuron, and the cell’s current state, characterized by the full set of conductances $\bar{g}_{\mathrm{ion}}\in\mathbb{R}^{N}$. The computation of these references is achieved through DICs, which encapsulate this complexity.
DICs consist of three voltage-dependent conductances that separate according to timescales: one fast, one slow, and one ultraslow, denoted as $g_{\mathrm{f}}(V)$, $g_{\mathrm{s}}(V)$, and $g_{\mathrm{u}}(V)$, which can be computed as linear functions of the maximal conductance vector $\bar{g}_{\mathrm{ion}}\in\mathbb{R}^{N}$ of an $N$-channel conductance-based model at each voltage level $V$:
$[g_{\mathrm{f}}(V);g_{\mathrm{s}}(V);g_{\mathrm{u}}(V)]=f_{\mathrm{DIC}}(V)=S(V)\cdot\bar{g}_{\mathrm{ion}}\,,$
where $S(V)\in\mathbb{R}^{3\times N}$ is a sensitivity matrix that can be built by: $S_{i\,j}(V)=-\left(w_{i\,j}\cdot\frac{\partial\dot{V}}{\partial X_{j}}\frac{\partial X_{j,\infty}}{\partial V}\right)/g_{\mathrm{leak}}$, where $i$ denotes the timescale, $X_{j}$ are gating variables of the j-th channel of the considered model and $w_{i\,j}$ is a timescale-dependent weight which is computed as the logarithmic distance of the time constant of $X_{j}$ and the timescale $i$ (Drion et al., 2015a, ). While the complete curve of the DICs may be of interest, only its value at the threshold voltage $V_{\mathrm{th}}$ is used, as the values and signs of the DICs at $V_{\mathrm{th}}$ reliably determine the firing pattern Drion et al., 2015a . Thus, the following linear system $f_{\mathrm{DIC}}(V_{\mathrm{th}})=S(V_{\mathrm{th}})\cdot\bar{g}_{\mathrm{ion}}$ makes the link between ion channel conductances and neuronal activity.
COMPETING INTEREST
The authors declare no competing interests.
FUNDING
Arthur Fyon is a Research Fellow of the “Fonds de la Recherche Scientifique - FNRS”, supported by grant ASP-REN40024838. This work was supported by the Belgian Government through the Federal Public Service (FPS) Policy and Support (BOSA) grant NEMODEI (NEuroMOrphic Design of Embodied Intelligence).
AUTHOR CONTRIBUTIONS STATEMENT
A.F. and G.D. conceived and designed the research. A.F. and G.D. developed the model code. A.F. performed the simulations and analyzed the data. A.F. and G.D. interpreted the results. A.F. prepared figures and drafted the manuscript. A.F. and G.D. edited and revised the manuscript. A.F. and G.D. approved the final version of the manuscript.
DATA AVAILABILITY
All code and data can be found on the first author’s GitHub (https://github.com/arthur-fyon/BIOCONTROL_2025).

### v2

@@
-Neuromodulation and homeostasis: complementary mechanisms for robust neural function
+Controlled neuromodulation naturally interacts with homeostasis to produce robust and modulable neuronal function
 Abstract
-Neurons depend on two interdependent mechanisms—homeostasis and neuromodulation—to maintain robust and adaptable functionality.
-Homeostasis stabilizes neuronal activity by adjusting ionic conductances, whereas neuromodulation dynamically modifies ionic properties in response to external signals.
+Neurons rely on two interdependent mechanisms – homeostasis and neuromodulation – to maintain robust and adaptable functionality.
+Homeostasis stabilizes neuronal activity by adjusting ionic conductances, whereas neuromodulation dynamically modifies ionic properties in response to external signals carried by neuromodulators.
 Combining these mechanisms in conductance-based models often produces unreliable outcomes, particularly when sharp neuromodulation interferes with homeostatic tuning.
 This study explores how a biologically inspired neuromodulation controller can harmonize with homeostasis to ensure reliable neuronal function.
-Using computational models of stomatogastric ganglion and dopaminergic neurons, we demonstrate that controlled neuromodulation preserves neuronal firing patterns while maintaining intracellular calcium levels.
+Using computational models of stomatogastric ganglion and dopaminergic neurons, we demonstrate that controlled neuromodulation preserves neuronal firing patterns while homeostasis simultaneously maintains target intracellular calcium levels.
 Unlike sharp neuromodulation, the neuromodulation controller integrates activity-dependent feedback through mechanisms mimicking G-protein-coupled receptor cascades.
@@
 Maximizing neuronal degeneracy enhances the likelihood of such intersections, enabling robust modulation and compensation for channel blockades.
-We further show that this controller pairing extends to network-level activity, reliably modulating central pattern generators in crustaceans.
-These findings suggest that targeting neuromodulation pathways—rather than ion channels directly—may offer safer pharmacological strategies to manage neuronal dysfunctions.
+We further show that this controller pairing extends to network-level activity, reliably modulating central pattern generators rhythmic activity.
 This study highlights the complementary roles of homeostasis and neuromodulation, proposing a unified control framework for maintaining robust and adaptive neural activity under physiological and pathological conditions.
-SIGNIFICANCE STATEMENT
+Author summary
 Neurons must maintain stable activity while adapting to changing demands, relying on two mechanisms: homeostasis, which stabilizes activity, and neuromodulation, which adjusts behavior to external signals.
 These mechanisms often interact, but their improper coordination can lead to dysfunction.
-This study uses computational models to show how controlled neuromodulation - mimicking biological feedback - can be harmonized with homeostasis to ensure robust neuronal function.
-By improving neuronal degeneracy, the system compensates for channel blockades and maintains critical activity patterns for both single neurons and networks.
+This study uses computational models to show how controlled neuromodulation – mimicking biological feedback – can be harmonized with controlled homeostasis to ensure robust neuronal function.
+The system compensates for channel blockades and maintains critical activity patterns for both single neurons and networks even when faced with neuronal parameter variability.
 These insights not only deepen our understanding of neural robustness, but also suggest safer pharmacological strategies targeting neuromodulatory pathways, offering potential breakthroughs in treating neurological disorders without disrupting essential cellular mechanisms.
 Introduction
-Brain activity is continuously shaped by neuromodulators and neuropeptides, including dopamine, serotonin, and histamine (Bargmann and Marder,, 2013; Marder et al.,, 2014).
-Neuromodulators dynamically influence single-neuron activity, input-output properties, and synaptic connection strength and dynamics, enabling neuronal networks to adapt to changing needs, contexts, and environments (Marder and Calabrese,, 1996; Marder and Bucher,, 2001; McCormick et al.,, 2020).
-To achieve this modulation, neuromodulators dynamically reshape the density, dynamics, and kinetics of many single cell transmembrane proteins, making whole brain functional signaling strongly dependent on the robustness and reliability of neuromodulation mechanisms at the molecular and cellular levels (Nadim and Bucher,, 2014; Hooper and Marder,, 1984).
-While experimental studies underscore the ubiquity of neuromodulation across nervous systems, the fundamental principles underlying this mechanism remain incompletely understood.
-As such, studying neuromodulation is a very arduous task, both experimentally and computationally, which explains our limited understanding of its underlying mechanisms.
-Neuromodulation coexists with cellular homeostasis, a mechanism that gradually adjusts neuronal membrane properties to maintain a target activity level (Desai,, 2003; Marder and Goaillard,, 2006; O’Leary et al.,, 2013).
-Neuronal cells are dependent on homeostatic mechanisms to maintain optimal functionality throughout the extended useful life of mammals (Marder,, 2011).
-Despite continuous turnover of transmembrane proteins, such as ion channels and receptors, occurring on varying temporal scales ranging from hours to weeks (O’Leary and Wyllie,, 2011), the robustness and excitability of neurons must remain undisturbed (Marder and Prinz,, 2002).
-By monitoring only intracellular calcium levels, the neuron orchestrates the regulation of all its ion channel conductances through homeostatic self-tuning rules (O’Leary et al.,, 2010; Pratt and Aizenman,, 2007).
-This raises the question of how two different mechanisms acting on the same targets (ion channels), each potentially having divergent goals, can co-exist and complement each other harmoniously (Marder et al.,, 2014; Marder and Prinz,, 2002).
-We propose to study this interaction between neuromodulation and homeostasis computationally, using conductance-based models.
-Conductance-based models provide an electrical representation of the neuronal membrane, where the membrane is modeled as an RC circuit, while the ion channels are modeled as non-linear conductances (Izhikevich,, 2007).
-These models are the closest equivalent of biological neurons, thanks to their ability to model any ion channel from experimental data following a methodology first proposed by Hodgkin and Huxley (Hodgkin and Huxley,, 1952).
-There exists a myriad of different conductance-based models that model various types of neuronal cells (Liu et al.,, 1998; Qian et al.,, 2014; Plant,, 1981; Wechselberger et al.,, 2006; Drion et al.,, 2011; Katori et al.,, 2010; Megwa et al.,, 2023).
-On the one hand, a homeostatic controller has been developed by O’Leary et al., (2014) and can be applied on top of any conductance-based model.
-In this controller, the neuron can adjust all ionic conductances up or down depending on its average intracellular calcium level.
-If the calcium level is lower than a reference value, ionic conductances are increased following homeostatic tuning rules and vice versa.
-Such rules force the tuning of ionic conductances along a single axis, that is, by preserving conductance ratios or correlations.
-On the other hand, Fyon et al., (2023) has developed a reliable neuromodulation controller and can also be applied to any conductance-based model.
-This controller can robustly tune the external activity of neurons, which means that neurons with various sets of ion channel properties that lead to similar firing activities (a property called degeneracy (Drion et al., 2015b, )) can be neuromodulated using the same control action.
-In this study, we first examine the interaction between homeostasis and sharp neuromodulation, emphasizing that improperly combining the homeostatic controller with neuromodulation may result in unreliable outcomes.
-Subsequently, we expand our analysis to explore the interaction between the homeostatic controller and a neuromodulation controller, both at the single-cell level and in rhythmic networks.
-This approach enables the development of a combined controller that integrates the strengths of both individual controllers, facilitating robust modulation and preservation of neuronal function.
+Brain activity is continuously shaped by neuromodulators and neuropeptides, including dopamine, serotonin, and histamine [2, 30].
+Neuromodulators dynamically influence single-neuron activity, input-output properties, and synaptic strength, enabling neuronal networks to adapt to changing needs, contexts, and environments [28, 27, 33].
+This modulation occurs through the regulation of transmembrane proteins, such as ion channels and receptors, which alters neuronal excitability and synaptic dynamics [35, 20].
+While experimental studies highlight the ubiquity of neuromodulation, the fundamental principles governing its effects remain incompletely understood, making computational approaches crucial for unraveling its mechanisms.
+Neuromodulation coexists with homeostatic plasticity, a process that gradually adjusts neuronal membrane properties to maintain a target activity level [8, 29, 37, 7, 50, 49, 51].
+Neuronal cells are dependent on homeostatic mechanisms to maintain optimal functionality throughout the extended useful life of mammals [32].
+Despite continuous turnover of transmembrane proteins, such as ion channels and receptors, occurring on varying temporal scales ranging from hours to weeks [39], the robustness and excitability of neurons must remain undisturbed [31].
+Among homeostatic mechanisms, intrinsic homeostatic plasticity adjusts ion channel conductances in response to intracellular calcium levels [36, 41].
+However, since both neuromodulation and homeostasis act on the same targets (ion channels) and may have competing objectives, understanding their interaction is essential [30, 31].
+We propose to study the interaction between neuromodulation and homeostasis using conductance-based models.
+These models provide a biophysically accurate representation of neuronal dynamics by describing the electrical properties of the neuronal membrane.
+The membrane is modeled as an RC circuit, while ion channels are represented as nonlinear conductances that evolve according to experimentally derived activation and inactivation kinetics [22].
+Conductance-based models remain the most faithful mathematical representations of biological neurons, as they allow for the direct incorporation of ion channel dynamics extracted from experimental data, following the methodology first introduced by Hodgkin and Huxley [19].
+Over the years, numerous conductance-based models have been developed to describe a variety of neuronal cell types across different species and brain regions [25, 43, 40, 52, 12, 23, 34].
+This study focuses on a conductance-based model of stomatogastric ganglion (STG) neurons [25].
+STG neurons are key components of rhythmic circuits that generate motor patterns driving stomach muscle contractions in crustaceans.
+This model was chosen for its ability to reproduce diverse firing patterns – ranging from tonic spiking to bursting – and its rich set of ion channels, which enable degeneracy.
+Specifically, the model includes classical fast sodium and delayed rectifier potassium channels, A-type and calcium-activated potassium channels, T-type and slow calcium channels, as well as hyperpolarization-activated H-channels.
+We model intrinsic homeostasis by using the homeostatic controller proposed by [38], which adjusts all ionic conductances up or down based on the neuron average intracellular calcium level.
+If the calcium level is lower than a target reference value, the controller increases ionic conductances according to homeostatic tuning rules, and conversely, decreases them if the calcium level is above the target.
+Importantly, these rules ensure that conductances are tuned along a single axis, preserving their relative ratios and maintaining correlations between them.
+This self-tuning mechanism allows neurons to counteract perturbations while preserving stable activity.
+Neuromodulation is modeled using two different approaches.
+On the one hand, we model neuromodulation as an acute change in target conductance values leading to a desired neuromodulated state in degenerate populations, which we call ”sharp neuromodulation”.
+On the other hand, we implement the neuromodulation controller developed by [14], which is designed to enable robust modulation of neuronal activity.
+We call this approach ”controlled neuromodulation”.
+One of its key properties is that it accounts for neuronal degeneracy, meaning that neurons with different ion channel compositions but similar activity patterns respond in a consistent manner to neuromodulation [15].
+This feature is particularly important because biological neuronal populations exhibit significant variability in their underlying conductance parameters while maintaining functional stability.
+By leveraging degeneracy, the neuromodulation controller can apply the same control signal across a population of neurons with diverse conductance profiles, ensuring reliable and predictable neuromodulation.
+We first investigate the interaction between intrinsic homeostatic plasticity and neuromodulation at the single-neuron level and show that combining intrinsic homeostasis with either sharp or controlled neuromodulation leads to vastly different outcomes.
+Secondly, we provide a mechanistic explanation for these different outcomes.
+We then extend our analysis to rhythmic neuronal networks, where neuromodulation plays a critical role in adjusting oscillatory patterns.
+By carefully integrating homeostatic and neuromodulatory control mechanisms, we aim to develop a unified framework that combines the strengths of both approaches.
+This integrated controller enables robust modulation of neuronal activity while ensuring long-term stability, shedding light on how neuromodulation and homeostasis coexist in biological systems.
 Results
-Current model of homeostasis can lead to pathological excitability in the presence of sharp neuromodulation
-This study employs a stomatogastric ganglion (STG) neuron conductance model (Liu et al.,, 1998) to explore the interaction between homeostasis and neuromodulation.
-These neurons contribute to rhythmic circuits that generate signals triggering stomach muscle contractions in crustaceans.
-This model was chosen for its ability to produce diverse firing patterns—ranging from tonic spiking to bursting, including single-spike bursting—and its rich combination of ion channels enabling degeneracy.
-The model includes classical fast sodium and delayed rectified potassium channels, A-type and calcium-activated potassium channels, T-type and slow calcium channels as well as hyperpolarization-activated H-channels.
-We first study the interaction between state-of-the-art model homeostasis (O’Leary et al.,, 2014) and sharp neuromodulation (Fyon et al.,, 2024).
-We define sharp neuromodulation as a direct change in a subset of ion channel maximal conductances, which results in functional modulation.
-In the context of this work, we will focus on a robust transition from tonic spiking to bursting in degenerate neurons, i.e.
-neurons exhibiting similar firing activities from different underlying channel conductances.
-First, we sharply neuromodulated a set of degenerate neurons from tonic spiking to bursting by directly modifying the maximal conductances of A-type potassium $\bar{g}_{A}$ and slow calcium $\bar{g}_{CaS}$, then allowed homeostatic compensation to take its course.
-Three example traces with a time evolution of maximal conductances are provided in Fig.
-1, left, whereas Fig.
-1, right show trajectories in the $\bar{g}_{\mathrm{CaS}}$-$\bar{g}_{\mathrm{A}}$ plane.
-As can be seen in the figure, although the initial tonic firing and the sharply neuromodulated bursting activities are very similar in all three neurons, homeostatic compensation has very variable and unreliable effects.
-Depending on the initial conductance state, neuromodulation shifts the neuron to a new state where homeostatic compensation can lead to pathological (Fig.
-1A), functionally preserved (Fig.
-1B), or unphysiological (Fig.
-1C) behaviors.
-This interesting result highlights an important property of neuronal degeneracy: although distinct channel subsets can produce similar neuronal behaviors, neurons may respond differently to perturbations, such as sharp neuromodulation.
-To better understand the reasons behind this unreliability, both mechanisms must be discussed in detail.
-In the homeostatic model O’Leary et al., (2014), the cell senses its own intracellular calcium level and averages it over time.
-If the mean intracellular calcium falls below a predefined target value, RNA transcription and ion channel protein translation are triggered to increase all ionic conductance values, and vice versa.
-Specifically, increasing calcium ion channel conductances leads to more calcium ions entering the cell.
-This implies that homeostasis induces changes in all conductance values (Fig.
-1 left), which leads to a linear motion in the conductance plane (Fig.
-1 right).
-This phenomenon happens on a relatively slow timescale, which can last up to days (O’Leary et al.,, 2010).
-Notably, the RNA transcription time constants, denoted by $\tau_{m}$, vary across different ion channels and cell types.
-Steady-state mathematical analysis has shown that these differing time constants tune correlations in ion channel conductances.
-For two ion channels $i$ and $j$, steady-state analysis yields $\bar{g}_{i}/\bar{g}_{j}\approx\tau_{m_{j}}/\tau_{m_{i}}$.
-This indicates that homeostatic tuning rules maintain ratios of ionic conductances and their correlations while adjusting their values in response to intracellular calcium levels.
-In the ionic conductance space, these rules allow neurons to navigate along a one-dimensional line, where the slope of this line in two dimensions is determined by the ratios of mRNA transcription time constants.
-This phenomenon, known as homogeneous scaling, involves proportionally increasing or decreasing all ionic conductances by the same multiplicative factor.
-On the other hand, sharp neuromodulation is implemented by transiently applying the algorithm of Fyon et al., (2024).
-This neuromodulation also produces changes in conductance values (Fig.
-1 left), resulting in motion within the conductance plane in a direction different from homogeneous scaling (Fig.
-1 right) to reach a new firing activity.
-However, neuromodulation only affects a subset of ion channels, not all (Marder,, 2012).
-Furthermore, the timescale of neuromodulation is much faster than that of homeostasis, lasting for a maximum of a few minutes (Marder et al.,, 2014).
-In addition, neuromodulation tunes correlations between neuromodulated ion channels, rather than preserving them (Fyon et al.,, 2024).
-Coupling homeostasis and neuromodulation in this way is problematic because homeostasis preserves correlations and regulates calcium, whereas neuromodulation adjusts correlations to modulate neuronal function.
-In Fig.
-1, sharp neuromodulation is applied to three different STG models with distinct initial ionic conductances.
-Initially, all neurons exhibited tonic firing due to homeostatic regulation, with RNA transcription time constants selected to produce tonic spiking neurons.
-Neuromodulation was then applied instantaneously, leading to a transition from tonic spiking to bursting in all cases (Fig.
-1).
-These transitions drastically increase intracellular calcium concentration.
-From the perspective of homeostatic tuning rules, this constitutes a disturbance, as it alters the conductance ratios between $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$.
-Although this disturbance is necessary for adaptation and functional modulation, the outcome of homeostasis in response to this biological disturbance is unpredictable.
-Indeed, as the homeostatic controller was initially designed to produce tonic spiking neurons, it tends to preserve the predefined correlations that produce such behavior.
-However, neuromodulation changes the firing activity, leading to a mismatch between actual conductance values and the encrypted conductance ratios.
-In Fig.
-1A, homeostatic compensation results in pathological behavior, characterized by irregular bursting with a hyperpolarized plateau, failing to achieve either bursting or tonic spiking.
-In Fig.
-1B, the modulated behavior (bursting) is maintained after homeostatic compensation.
-In Fig.
-1C, homeostatic compensation leads to unphysiological values of $\bar{g}_{\mathrm{A}}$, tending towards negative values.
-Even when $\bar{g}_{\mathrm{A}}$ is constrained to be non-negative, the resulting electrical behavior is pathological, failing to exhibit either tonic spiking or bursting.
-This toy experiment highlights that homeostasis and sharp neuromodulation can interfere pathologically.
-Such behavior was predicted in O’Leary et al., (2014), where the authors demonstrate that their homeostasis model can lead to various outcomes in response to disturbances, such as channel deletions.
-Initially, the deletion itself may directly affect neuronal function, either preserving it or causing its loss.
-Subsequently, homeostatic compensation can either preserve or restore function, but it may also lead to pathological behaviors, such as functional deletion.
-This result suggests that any disturbance applied to the conductance ratios—even physiological ones such as neuromodulation—can lead to problematic behaviors.
-Figure 1: Current model of homeostasis can lead to pathological excitability in the presence of sharp neuromodulation.
-A.
-Time evolution of all conductances in the STG model, shown on a logarithmic scale, when homeostasis is applied with sharp neuromodulation (instantaneous changes in $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$ within the red region) (left).
-Corresponding trajectory in the modulated conductance space (right).
-The white circle represents the initial condition (tonic spiking), the red diamond marks the neuromodulated state (bursting) immediately after the sharp modulation, and the green triangle denotes the final state after homeostasis has converged.
-In this case, homeostatic compensation is pathological, resulting in irregular bursting.
-B.
-Same as in panel A, but with different degenerate initial conditions.
-Here, homeostatic compensation preserves the neuromodulated function.
-C.
-Same as in panel A, with another set of degenerate initial conditions.
-In this case, homeostatic compensation results in negative values of $\bar{g}_{\mathrm{A}}$ (saturated at 0), leading to pathological behavior characterized by one-spike bursting.
-Controlled neuromodulation and homeostasis naturally lead to robust neuronal function
-In biological systems, homeostasis and neuromodulation collaborate to ensure robust and modular neural function.
-These mechanisms operate on different timescales: neuromodulation acts within seconds or minutes (Werner and Mitterauer,, 2014), while homeostatic compensation may take hours or even days (Marder and Prinz,, 2002).
-However, in computational models, the goals of neuromodulation and homeostatic controllers often seem discordant.
-Neuromodulation typically adjusts conductance ratios, while homeostasis seeks to maintain these ratios and guide the neuron through homogeneous scaling.
-Consequently, applying sharp neuromodulation to any homeostatic conductance-based model can lead to instability, as neuromodulation disrupts the references and control actions of the homeostasis controller (Fig.
-2A).
-To model the actions of both mechanisms within a single cell, we include a more advanced neuromodulation model, referred to as the neuromodulation controller, as developed in Fyon et al., (2023).
-This controller can be coupled with any conductance-based model, allowing modulation of its neuronal function, such as its firing pattern, to mimic biological neuromodulation.
-We call this type of neuromodulation controlled neuromodulation, in contrast to the sharp neuromodulation discussed above.
-The neuromodulation controller adjusts target channel conductances, providing continuous and consistent modulation rather than instantaneous, rigid changes.
-Fig.
-2B illustrates the results of transitions from tonic spiking to bursting when applied to a degenerate population of STG neuron models.
-Initially, all neurons exhibit tonic firing mode, and homeostasis adjusts the conductances to achieve a predefined calcium level along a first tonic spiking homogeneous scaling line.
-Midway through the simulation, controlled neuromodulation is applied to switch all neurons to bursting mode, which now occurs through a controlled adjustment of $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$.
-Consequently, intracellular calcium levels rise sharply, and conductance ratios involving one or both neuromodulated ionic conductances are adjusted.
-Finally, with the reshaped RNA time constants, the homeostasis controller adjusts all conductances to restore an appropriate calcium level, while controlled neuromodulation preserves neuromodulated conductance ratios along a second bursting homogeneous scaling line.
-Ultimately, the neuronal population reaches a new state where the neuromodulated function is preserved and intracellular calcium levels are back to their target value.
-Simulations were also performed on a midbrain dopaminergic neuron model (adapted from Qian et al., (2014)) to demonstrate the generality of our results (Fig.
-2C).
-As with the previous simulations, a transition from tonic spiking (pacemaking) to strong bursting was induced midway through by modulating the N-type and L-type calcium conductances ($\bar{g}_{\mathrm{CaN}}$ and $\bar{g}_{\mathrm{CaL}}$) using controlled neuromodulation.
-Homeostatic compensation was continuously applied to maintain proper calcium levels, resulting in a reliable and robust outcome despite initial degeneracy within the population.
-Figure 2: Controlled neuromodulation and homeostasis naturally lead to robust neuronal function.
+Interference between homeostasis and sharp neuromodulation leads to pathological outcomes
+We first examine the interplay between a state-of-the-art model of cellular homeostasis [38] and sharp neuromodulation [13].
+Here, sharp neuromodulation is defined as a direct, externally driven modification of a subset of ion channel maximal conductances, resulting in a change in neuronal spiking behavior.
+We specifically focus on a robust transition from tonic spiking to bursting in degenerate neuronal populations, i.e.
+populations composed of neurons that exhibit similar firing behaviors despite differences in underlying conductance profiles.
+To explore this interaction, we applied sharp neuromodulation to a population of degenerate neurons generated using the dataset from [13] (Fig 1).
+Sharp neuromodulation consisted of directly modifying the maximal conductances of A-type potassium ($\bar{g}_{\mathrm{A}}$) and slow calcium ($\bar{g}_{\mathrm{CaS}}$) channels to transition neurons from tonic spiking to bursting.
+We then allowed homeostatic compensation to unfold.
+Since bursting increases overall neuronal activity, intracellular calcium levels rise, triggering a decrease in all conductances according to homeostatic tuning rules.
+Three representative conductance trajectories ($N=200$) are shown in Fig 1 (top), while Fig 1 (bottom) illustrates calcium dynamics across the population (note that to facilitate observation of both mechanisms within a single trace, the homeostatic controller was accelerated compared to its original implementation).
+Despite initially similar firing patterns and a similar outcome of sharp neuromodulation, homeostatic compensation produced highly variable and often unreliable outcomes in terms of modulated activity, while still regulating intracellular calcium.
+Depending on the initial conductance state, neuromodulation led to new states where homeostasis could either maintain functional activity, result in pathological behavior, or drive neurons into unphysiological regimes.
+This finding underscores a fundamental property of degeneracy: neurons with different underlying conductances can exhibit similar activity yet respond differently to perturbations, such as neuromodulation.
+Fig 1: Interference between homeostasis and sharp neuromodulation leads to pathological outcomes.
+Time evolution of all conductances in the STG model, displayed on a logarithmic scale, during homeostasis with sharp neuromodulation (instantaneous changes in $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$ starting at the dashed blue line) for a degenerate population of 200 models (top).
+The corresponding mean intracellular calcium concentration over time (bottom) shows the target value (red dash-dotted line) is maintained both before and after sharp neuromodulation.
+To understand the source of this unreliability, it is necessary to examine both mechanisms in more detail.
+The homeostatic model [38] relies on intracellular calcium as a feedback signal.
+The neuron continuously averages its intracellular calcium level over time and adjusts conductances accordingly: if calcium falls below a target level (red dash-dotted line in Fig 1 bottom), ion channel transcription and translation increase, raising all conductance values, and vice versa.
+This phenomenon happens on a relatively slow timescale, which can last up to days [36].
+In contrast, sharp neuromodulation, implemented using the method from [13], directly modifies a subset of ion channel conductances, which upsets the homeostasis setpoint, defined by the intracellular calcium target level.
+As homeostasis kicks in to recover its calcium set point, there are no guarantees that the neuromodulated function is maintained as ion channel conductances are modified, which often leads to pathological or unphysiological outcomes.
+This type of interference was anticipated in [38], where the authors demonstrated that homeostatic regulation can produce diverse responses to perturbations such as ion channel deletions.
+Depending on the initial conditions, a deletion may either preserve function, disrupt it, or lead to pathological compensation.
+Our results extend this idea by showing that even physiological disturbances, such as neuromodulation, can induce similar unpredictability when coupled with homeostatic regulation.
+Controlled neuromodulation and homeostasis naturally lead to robust and modulable neuronal function
+In biological systems, neuromodulation occurs through an intracellular cascade that indirectly links changes in neuromodulation concentration and changes in targeted conductance values using an intermediate signal such as e.g.
+cyclic amp.
+We have recently shown that this indirect action of neuromodulation on conductance values could be viewed as an intracellular controller (Fig 2A, top), which largely improves reliability of neuromodulatory action on highly degenerate populations [14].
+We call this mechanism ”controlled neuromodulation” to contrast with the sharp neuromodulation described above.
+Unlike sharp neuromodulation, which imposes instantaneous conductance changes, the neuromodulation controller provides continuous and adaptive modulation, ensuring stable functional transitions.
+Fig 2: Controlled neuromodulation and homeostasis naturally lead to robust and modulable neuronal function.
 A.
@@
 B.
-Time evolution of all conductances in the STG model, displayed on a logarithmic scale, during homeostasis with controlled neuromodulation (controlled changes in $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$ starting at the dashed blue line) for a degenerate population of 200 models (left).
+Time evolution of all conductances in the STG model, displayed on a logarithmic scale, during homeostasis with controlled neuromodulation (controlled changes in $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$) for a degenerate population of 200 models (left).
 The corresponding mean intracellular calcium concentration over time (right) shows the target value (red dash-dotted line) is maintained both before and after controlled neuromodulation.
-C.
-Same as panel B, but for a DA model where controlled neuromodulation affects $\bar{g}_{\mathrm{CaN}}$ and $\bar{g}_{\mathrm{CaL}}$.
-Why do controlled neuromodulation and homeostasis lead to robust neural function?
-Fig.
-3A shows the trajectories of the simulations from Fig.
-2B in the plane of modulated conductances, $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$, during tonic spiking homeostatic compensation for both sharp and controlled neuromodulation.
-Initially, the modulated conductances of every neuron in the population grow along the homogeneous scaling direction in tonic spiking mode, as the calcium levels are below the threshold value.
-This homogeneous scaling corresponds to a line intersecting the origin, with variability due to the degeneracy in conductance ratios within the population.
-During neuromodulation, all conductances are adjusted to tune their ratios and achieve bursting (Fig.
-3B).
-Importantly, all neuromodulatory actions occur along the same direction, but with varying magnitudes depending on the state of the neuron (Fyon et al.,, 2024).
-As a result, calcium levels increase drastically for each neuron in the population.
-Once the neuromodulatory action ends, the outcome depends on the type of neuromodulation applied (Fig.
-3C).
-In the case of sharp neuromodulation, during the homeostatic action in the bursting phase, the homeostatic controller acts alone and tends to maintain the same conductance ratios as in tonic spiking (Fig.
-3C left).
-This leads to a reduction in calcium levels, but, as seen in Fig.
-1, its effects are unreliable and undesirable.
-Specifically, for most neurons in the population, the homeostatic controller reduces $\bar{g}_{\mathrm{A}}$ towards negative values (saturated at 0 to avoid model instability) to bring $\bar{g}_{\mathrm{CaS}}$ into an appropriate range.
-This regulation causes the neuron to move away from the robust bursting behavior, resulting in heterogeneous changes in the activity of the population.
-In the case of controlled neuromodulation, the homeostatic and neuromodulation controllers work together to maintain the new conductance ratios corresponding to bursting (Fig.
-3C right).
-This results in a new homogeneous scaling following these updated conductance ratios in bursting, which translates into trajectories that move toward the origin of the axes, as would be expected in tonic spiking.
-Homogeneous scaling during bursting differs across the population due to initial degeneracy.
-The difference between the two homogeneous scalings (tonic spiking and bursting) is driven by the neuromodulatory action, ensuring that the neuromodulated behavior (bursting) is not lost.
-Figure 3: Why do controlled neuromodulation and homeostasis lead to robust neural function?
+Using the same protocol as for sharp neuromodulation, we now study the interaction between controlled neuromodulation and homeostasis.
+Fig 2B illustrates the transition from tonic spiking to bursting in a degenerate population of STG neuron models, identical to those in Fig 1, but when using controlled neuromodulation instead of sharp neuromodulation.
+Initially, all neurons exhibit tonic firing, with homeostasis tuning conductances along a tonic spiking homogeneous scaling line to maintain a target calcium level.
+Midway through the simulation, controlled neuromodulation is applied, gradually adjusting $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$ to induce bursting.
+As in the case of sharp neuromodulation, this transition increases intracellular calcium levels, leading to an adaptive shift in conductance values.
+Homeostatic regulation then readjusts overall conductance magnitudes, but here the interaction between homeostasis and controller neuromodulation ensures that the newly established neuromodulated state is maintained.
+Ultimately, the population reaches a stable configuration where both the neuromodulated function and intracellular calcium homeostasis are fulfilled.
+To demonstrate the generality of this approach, we applied the same framework to a midbrain dopaminergic neuron model adapted from [43] (see Supporting Information, S1 Appendix).
+As in the STG model, a transition from tonic spiking (pacemaking) to strong bursting was induced by modulating N-type and L-type calcium conductances ($\bar{g}_{\mathrm{CaN}}$ and $\bar{g}_{\mathrm{CaL}}$) through controlled neuromodulation.
+Homeostatic compensation continuously regulated calcium levels, ensuring a reliable and robust outcome despite initial conductance degeneracy within the neuronal population.
+Controlled neuromodulation tunes the direction of homeostasis-driven changes in conductance values to ensure stable neuromodulated function during homeostatic regulation
+To study the mechanisms underlying the constructive interaction between controlled neuromodulation and homeostasis, we plot the conductance trajectories of the simulations from Fig 1 in the plane of modulated conductances, $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$, during tonic spiking homeostatic compensation for both sharp and controlled neuromodulation (Fig 3A).
+Initially, as calcium levels remain below the threshold, homeostasis increases the modulated conductances along a homogeneous scaling direction characteristic of tonic spiking.
+This direction corresponds to a line intersecting the origin, with variability due to conductance degeneracy within the population, and its slope is determined by the relative time constants of mRNA transcription in the homeostasis model, with steady-state analysis predicting $\bar{g}_{i}/\bar{g}_{j}\approx\tau_{m_{j}}/\tau_{m_{i}}$ [38].
+When neuromodulation is applied, all conductances are adjusted to reshape their ratios and induce bursting (Fig 3B).
+Importantly, neuromodulatory actions follow the same direction across neurons but vary in magnitude depending on individual neuronal states [13].
+This modulation leads to a sharp increase in intracellular calcium levels.
+Once neuromodulation ceases, the long-term outcome depends on the type of neuromodulation applied (Fig 3C).
+Fig 3: Controlled neuromodulation tunes the direction of homeostasis-driven changes in conductance values to ensure stable neuromodulated function during homeostatic regulation.
 A.
-Trajectories of the STG population from Fig.
-2 in the modulated conductance space during homogeneous scaling in tonic spiking only.
+Trajectories of the STG population from Fig 1 in the modulated conductance space during homogeneous scaling in tonic spiking only.
 As expected from homeostasis, all neurons move in the same direction.
 White circles represent the initial conditions at the start of the step, and the green triangle marks the final conditions.
-B.
-Same as panel A, but during the brief period when neuromodulation is active.
-All trajectories are parallel to one another.
+B. Same as panel A, but during the brief period when neuromodulation is active. All trajectories are parallel to one another.
 C.
@@
 Sharp neuromodulation results in non-robust outcomes, whereas controlled neuromodulation preserves robust neural function.
-In the first section, when the homeostasis controller was combined with sharp neuromodulation, unreliable outcomes were observed.
-To explain this behavior, we use a schematic representation of the modulated conductance plane (Fig.
-4A).
-In this plane, different values of $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$ can lead to similar mean intracellular calcium concentration.
-These values can be highlighted using calcium isoclines (green dashed lines in Fig.
-4A).
-Likewise, different values of $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$ can lead to similar firing activity.
-These values can be highlighted using activity isoclines (blue dashed lines in Fig.
-4A).
-Calcium isoclines are parallel one to each other, whereas activity isoclines pass through the origin and rotate around this point.
-This behavior is explained by homeostatic tuning rules and homogeneous scaling, as different conductance ratios lead to different firing patterns.
-Initially, the neuron operates in the target spiking mode at the target calcium level, corresponding to the intersection of the spiking and the target calcium isoclines.
-When sharply neuromodulating the neuron, the modulated conductances change, and the neuron shifts to another state, now reaching a new target strong bursting isocline, which intersects a higher calcium isocline.
-Homeostasis then kicks in to reduce the calcium level to its target value.
-However, in the homeostatic controller, the “native” tonic spiking direction is encrypted, leading to changes in conductances along the spiking isocline.
-As a result, the neuron moves away from the bursting isocline, potentially disrupting the new function.
-This explains why sharp neuromodulation is incompatible with the homeostasis controller.
-To harmonize the actions of both controllers and achieve robust functional modulation, we leverage the significant difference in their timescales.
-Neuromodulation controls the modulated ionic conductances, while homeostatic tuning rules regulate all ionic conductances over a slower timescale to maintain an appropriate calcium level.
-In the modulated conductance plane (Fig.
-4B), the initial and neuromodulated states are the same as in Fig.
-4A.
-However, in the case of controlled neuromodulation, both homeostatic and neuromodulatory controllers are active.
-Specifically, when homeostasis reduces the calcium level and shifts the neuron away from the bursting isocline, the neuromodulation action counteracts this perturbation in function by bringing the neuron back to the bursting isocline, preserving its function.
-The neuromodulatory controller operates several orders of magnitude faster than the homeostatic controller.
-As a result, fast corrections from the former render perturbations from the latter infinitesimal, leading to changes in the conductance ratio along the target activity isocline (strong bursting in this example).
-Ultimately, the neuron stabilizes at the intersection of the target calcium and target activity isoclines, where both controllers meet their objective.
-Figure 4: Schematic explanation of why controlled neuromodulation integrates better with homeostasis compared to sharp neuromodulation.
+Under sharp neuromodulation, homeostasis operates alone in the bursting phase, following the scaling direction characteristic of its reference, i.e.
+unmodulated, state set by the relative time constants of mRNA transcription in the homeostasis model (Fig 3C, left).
+This results in a decrease in calcium levels, but as shown in Fig 1, the response is unreliable and often undesirable.
+Specifically, for most neurons, the homeostatic controller reduces $\bar{g}_{\mathrm{A}}$ toward negative values (saturated at 0 to prevent model instability) to compensate for $\bar{g}_{\mathrm{CaS}}$, driving the neuron away from robust bursting and leading to heterogeneous activity changes across the population.
+In contrast, controlled neuromodulation allows the homeostatic and neuromodulation controllers to act together, maintaining the new conductance ratios associated with bursting (Fig 3C, right).
+As a result, homeostatic regulation follows a new homogeneous scaling trajectory aligned with bursting conductance ratios.
+These trajectories converge toward the origin, as in tonic spiking, but with a distinct direction determined by neuromodulation.
+Due to initial conductance degeneracy, the exact scaling varies across the population.
+The shift between the two homogeneous scalings – tonic spiking and bursting – emerges from neuromodulation, ensuring that the desired neuromodulated function is maintained.
+The incompatibility of sharp neuromodulation with homeostasis can be further analyzed using a schematic representation of the modulated conductance plane (Fig 4A).
+In this plane, different combinations of $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$ can yield similar calcium levels, forming calcium isoclines (green dashed lines).
+Similarly, different conductance values can result in similar firing activity, forming activity isoclines (blue dashed lines).
+While calcium isoclines are parallel, activity isoclines radiate from the origin and rotate around it.
+This organization reflects homeostatic tuning rules: different conductance ratios correspond to distinct firing modes [38].
+Initially, the neuron operates at the intersection of a target spiking isocline and a target calcium isocline.
+When sharp neuromodulation is applied, conductances shift to a new state corresponding to strong bursting, moving the neuron to a higher calcium isocline.
+Homeostasis then acts to restore calcium levels, but because it retains the ”native” tonic spiking scaling direction, it alters conductances along the spiking isocline rather than preserving the newly induced bursting state.
+As a result, the neuron drifts away from the bursting isocline, potentially disrupting its function.
+This explains why sharp neuromodulation leads to unreliable outcomes when coupled with homeostasis.
+Fig 4: Schematic explanation of why controlled neuromodulation integrates better with homeostasis compared to sharp neuromodulation.
 A.
@@
 Because controlled neuromodulation operates on a much faster timescale than homeostasis, the neuron remains on the neuromodulated firing pattern isocline throughout.
+To reconcile homeostasis and neuromodulation and achieve robust functional modulation, we leverage their distinct timescales.
+Neuromodulation controls specific ionic conductances, while homeostasis adjusts all conductances over a slower timescale to regulate calcium levels.
+In the modulated conductance plane (Fig 4B), the neuron initially follows the same trajectory as in Fig 4A.
+However, under controlled neuromodulation, both controllers remain active.
+When homeostasis lowers calcium levels and shifts the neuron away from the bursting isocline, the neuromodulatory controller counteracts this deviation, bringing the neuron back toward the bursting isocline and preserving the desired function.
+Because the neuromodulatory controller operates orders of magnitude faster than the homeostatic controller, its rapid corrections render homeostatic perturbations negligible.
+This interaction constrains changes in conductance ratios along the target activity isocline (strong bursting in this case).
+Ultimately, the neuron stabilizes at the intersection of the target calcium and target activity isoclines, where both controllers achieve their objectives.
 Controlled neuromodulation and homeostasis ensure the preservation of function under physiologically recoverable disturbances
 By employing this tandem of controllers, neuronal function becomes robust to channel blockade, provided that the channel deletion is compensable.
-While homeostasis compensation alone might lead to unreliable recovery from channel blockade, incorporating controlled neuromodulation somehow stabilizes the neuron response to such blockade, even in the case of highly degenerate neuronal populations.
-Furthermore, if the channel blockade is physiologically recoverable—meaning the function of the blocked channel can be restored using other channels—there exists an optimal combination of neuronal feedback gains to preserve that function.
-In essence, if one channel is blocked, a specific concentration of particular neuromodulators can preserve the function.
-For example, in Fig.5, various channel blockades are applied to a degenerate population of STG conductance-based models equipped with the combined homeostasis/neuromodulation tandem of controllers.
+While homeostasis alone may lead to unreliable recovery, incorporating controlled neuromodulation stabilizes the neuronal response to blockade, even in highly degenerate neuronal populations.
+Furthermore, if the blocked channel function can be physiologically recovered – meaning other channels can substitute for its role – there exists an optimal combination of neuronal feedback gains that preserves the function.
+In essence, when a channel is blocked, a specific concentration of neuromodulators can maintain the function.
+For example, in Fig 5, various channel blockades are applied to a degenerate population of STG conductance-based models equipped with the combined homeostasis/neuromodulation tandem.
 The target behavior of the population is regular bursting.
-In Fig.5A, H-type channels are blocked.
-Here, the function is preserved even immediately after the blockade, without requiring compensation.
-Subsequently, the combined regulation of homeostasis and controlled neuromodulation maintains this function by restoring the neuron to an appropriate calcium level on a population-wide scale.
-In Fig.
-5B and C, T-type calcium and calcium-controlled potassium channels are blocked, respectively.
-In these cases, the blockade results mostly in the loss of bursting function.
-However, by allowing the homeostasis/neuromodulation pair of controllers to compensate, the function is largely recovered, albeit with irregular bursting patterns in some neurons.
-Additionally, intracellular calcium levels stabilize at their target values.
-This preservation of functions is feasible because the actions of H-type, T-type calcium, and calcium-controlled potassium channels are degenerate and can be compensated for using other channels present within the cell.
-The STG neuron model incorporates numerous channels operating on slow and ultraslow timescales, enabling the compensation for lost channels in these timescales.
-It is worth noting that only the loss of the calcium-controlled potassium channel has been compensated using different neuronal feedback gain values, i.e., different concentrations of various neuromodulators, as in control conditions.
-However, the blockade of fast transient sodium channels is irrecoverable (Fig.
-5D), as these channels are crucial for initiating spikes on a fast timescale.
-Once the blockade is applied, compensation cannot restore spikes because the action of sodium channels is non-degenerate and occurs on the fast timescale of the neuron.
-Similarly, achieving the appropriate intracellular calcium level becomes unattainable.
-A comparable outcome would result from the blockade of delayed rectified potassium channels, which are responsible for the downstroke of the spikes.
-Figure 5: Controlled neuromodulation and homeostasis ensure the preservation of function under physiologically recoverable disturbances.
+In Fig 5A, H-type channels are blocked.
+Here, the function remains intact even immediately after the blockade, without requiring compensation.
+The combined action of homeostasis and controlled neuromodulation then ensures that intracellular calcium levels are restored on a population-wide scale, maintaining the function.
+Fig 5: Controlled neuromodulation and homeostasis ensure the preservation of function under physiologically recoverable disturbances.
 A.
@@
 In this case, no neurons lose bursting due to the channel blockade or compensation.
-B.
-Same as panel A, but with a CaT channel blockade.
-The blockade causes half of the neurons to lose bursting; however, all neurons recover bursting with compensation.
-C.
-Same as panel A, but with a KCa channel blockade.
-The blockade causes more than half of the neurons to lose bursting, but nearly all recover bursting with compensation.
+B. Same as panel A, but with a CaT channel blockade. The blockade causes half of the neurons to lose bursting; however, all neurons recover bursting with compensation.
+C. Same as panel A, but with a KCa channel blockade. The blockade causes more than half of the neurons to lose bursting, but nearly all recover bursting with compensation.
 D.
@@
 The blockade causes all neurons to lose bursting, and compensation does not restore bursting due to the non-degenerate essential role of Na channels.
+In Fig 5B and C, T-type calcium and calcium-activated potassium channels are blocked, respectively.
+In these cases, the blockade largely disrupts bursting.
+However, allowing the homeostasis/neuromodulation controllers to compensate restores function in most neurons, albeit with some displaying irregular bursting.
+Additionally, intracellular calcium levels stabilize at their target values.
+This recovery is possible because the actions of H-type, T-type calcium, and calcium-activated potassium channels are degenerate, allowing their function to be compensated by other channels in the cell.
+The STG neuron model includes numerous channels operating on slow and ultraslow timescales, enabling compensation over these timescales.
+Notably, only the loss of the calcium-activated potassium channel has been compensated by adjusting neuronal feedback gains – equivalent to modifying neuromodulator concentrations – as in control conditions.
+In contrast, the blockade of fast transient sodium channels is irrecoverable (Fig 5D), as these channels are essential for spike initiation on a fast timescale.
+Once the blockade is applied, compensation cannot restore spiking because sodium channel function is non-degenerate and operates on a timescale that alternative channels cannot compensate for.
+Consequently, intracellular calcium levels also fail to reach their target values.
+A similar outcome would occur following the blockade of delayed rectifier potassium channels, which are responsible for the downstroke of spikes.
 Central pattern generator networks can be robustly modulated to orchestrate their rhythmic output
-The application of this tandem of controllers extends even to networks, particularly central pattern generators (CPGs).
-CPGs are self-organized neuronal circuits mainly composed of bursting neurons connected primarily through inhibitory connections, capable of generating rhythmic output without external input (Yuste et al.,, 2005; Guertin,, 2009).
-One extensively studied circuit is responsible for the pyloric and gastric mill rhythms in crustaceans, observed in the stomatogastric ganglion.
-This network exhibits two rhythms at different speeds: the fast pyloric rhythm and the slow gastric mill rhythm.
-These rhythmic outputs are vital for coordinating muscle contractions in the stomach for digestion, among other functions.
-A minimalist network model of such a bi-rhythmic circuit is investigated in Fig.
-6 (Gutierrez et al.,, 2013; Drion et al.,, 2019).
-It comprises two half-center oscillators operating at different speeds, connected via a central neuron.
-Half-center oscillators consist of two bursting neurons connected in mutual inhibition, enabling them to fire anti-phase bursts.
-In the network scheme, the purple and blue neurons constitute the fast half-center oscillator (representing the pyloric rhythm), while the green and yellow neurons form the slow half-center oscillator (representing the gastric mill rhythm).
-The red neuron serves as the central one, also in bursting mode.
-It has been demonstrated that the gastric mill rhythm requires a neuromodulatory input to be activated (Selverston et al.,, 2009).
-In this study, the gastric mill rhythm can be reliably induced by the pair of controllers, but not with homeostasis combined with sharp neuromodulation.
-Under control conditions (Fig.
-6A), i.e., without neuromodulatory input to the gastric mill rhythm, only the pyloric rhythm is active and the central neuron consistently bursts.
-In particular, variability in conductances is observed in this network (Fig.
-6A right).
-When neurons are equipped with homeostasis only, activation of sharp neuromodulation for the gastric mill rhythm leads to rapid loss of function, due to the unreliability of this combination (Fig.
-6B middle).
-However, when neurons undergo controlled neuromodulation, the gastric mill rhythm is successfully induced and sustained (Fig.
-6B right).
-Despite the degeneracy of neurons, characterized by significant variability in conductance ratios, the network achieves rhythmic output with appropriate intracellular calcium levels.
-This toy experiment demonstrates that such a tandem of controllers can be applied in computational models to modulate and maintain not only single cell activity but also network-level activity.
-Figure 6: Central pattern generator networks can be robustly modulated to orchestrate their rhythmic output.
+The application of this tandem of controllers extends beyond single neurons to networks, particularly central pattern generators (CPGs).
+CPGs are self-organized neuronal circuits composed primarily of bursting neurons connected through inhibitory synapses, capable of generating rhythmic output without external input [55, 17].
+A well-studied example is the circuit responsible for the pyloric and gastric mill rhythms in crustaceans, located in the stomatogastric ganglion.
+This network produces two distinct rhythms: the fast pyloric rhythm and the slow gastric mill rhythm.
+These rhythmic outputs are essential for coordinating muscle contractions in the stomach for digestion, among other functions.
+A minimal network model of such a bi-rhythmic circuit is investigated in Fig 6 [18, 11].
+It consists of two half-center oscillators operating at different speeds, connected via a central neuron.
+Half-center oscillators are formed by two bursting neurons connected through mutual inhibition, allowing them to fire anti-phase bursts.
+In this model, the purple and blue neurons constitute the fast half-center oscillator (representing the pyloric rhythm), while the green and yellow neurons form the slow half-center oscillator (representing the gastric mill rhythm).
+The red neuron acts as the central element, also in bursting mode.
+It has been shown that the gastric mill rhythm requires neuromodulatory input to be activated [47].
+Fig 6: Central pattern generator networks can be robustly modulated to orchestrate their rhythmic output.
 A.
-Control condition of the simplified pyloric/gastric mill rhythm network (left), inspired from Gutierrez et al., (2013).
+Control condition of the simplified pyloric/gastric mill rhythm network (left), inspired from [18].
 Without neuromodulation, the fast rhythm is active (top two traces, middle), while the slow rhythm is inactive (bottom two traces, middle).
@@
 With sharp neuromodulation, the slow rhythm fails due to its unreliability, whereas controlled neuromodulation sustains the slow rhythm.
+In this study, the tandem of controllers reliably induces the gastric mill rhythm, whereas homeostasis combined with sharp neuromodulation does not.
+Under control conditions (Fig 6A), i.e., in the absence of neuromodulatory input to the gastric mill rhythm, only the pyloric rhythm remains active, with the central neuron bursting continuously.
+Notably, conductance variability is observed within the network (Fig 6A right).
+When neurons are equipped with homeostasis alone, the activation of sharp neuromodulation for the gastric mill rhythm leads to a rapid loss of function due to the unreliability of this combination (Fig 6B middle).
+However, with controlled neuromodulation, the gastric mill rhythm is successfully induced and sustained (Fig 6B right).
+Despite the intrinsic degeneracy of the neurons, reflected in significant variability in conductance ratios, the network achieves stable rhythmic output while maintaining appropriate intracellular calcium levels.
+This toy experiment demonstrates that the tandem of homeostasis and controlled neuromodulation can be applied in computational models to modulate and sustain not only single-neuron activity but also network-level dynamics.
 Discussion
-The efficient and robust integration of homeostasis and neuromodulation in neurons remains an open question with several proposed hypotheses (Marder et al.,, 2014; Marder and Bucher,, 2005; Cunha,, 2001).
-The improper integration of these mechanisms—for example, combining sharp neuromodulation with a homeostatic controller—can lead to negative interactions.
-Such interactions may result in unreliable behaviors or numerical instabilities.
-In contrast, pairing a homeostatic controller with controlled neuromodulation ensures reliable neural function modulation while maintaining appropriate calcium levels.
-This approach draws inspiration from biology, where neuromodulation is known to be activity-dependent, as observed experimentally (Kramer and Levitan,, 1990; Raymond et al.,, 1992).
-Specifically, the neuromodulation controller models the role of G-protein-coupled receptors and second messengers, embedding feedback in the metabolic neuromodulation cascade (Marcus and Bruchas,, 2023; Scheler,, 2004).
-This reliable interaction can be extended to neural circuits, such as the gastric mill circuit in crabs, to produce robust network-level functionality (Marder et al.,, 2014).
+The efficient and robust integration of homeostasis and neuromodulation in neurons remains an open question, with several hypotheses proposed to explain their interaction [30, 6].
+Improper integration of these mechanisms – for instance, combining sharp neuromodulation with a homeostatic controller – can lead to negative interactions, resulting in unreliable behaviors or numerical instabilities.
+In contrast, pairing a homeostatic controller with controlled neuromodulation enables reliable modulation of neural function while maintaining appropriate calcium levels.
+This approach is biologically inspired, as neuromodulation is known to be activity-dependent, as observed experimentally [24, 45].
+Specifically, the neuromodulation controller models the role of G-protein-coupled receptors and second messengers, embedding feedback within the metabolic neuromodulation cascade [26, 46].
+This reliable interaction extends to neural circuits, such as the gastric mill circuit in crabs, where it contributes to robust network-level function [30].
 However, the success of this interaction critically depends on a point in the conductance space where both controllers converge.
 This point represents a neural activity pattern that supports neuromodulated firing while maintaining homeostatic calcium levels.
-As shown in Fig.
-4, this intersection of the strong bursting isocline (the target of controlled neuromodulation) and the target calcium level (the goal of homeostatic control) represents the convergence point.
-When such a point is absent—e.g., during sodium channel blockade (Fig.
-5)—neither neuromodulated activity nor calcium homeostasis is sustained.
-This may lead to pathological compensations, where one controller prioritizes calcium regulation at the cost of abandoning the desired firing pattern (Starmer et al.,, 1991; Grant et al.,, 1984; Anger et al.,, 2001).
-Furthermore, discontinuities in the neuromodulated firing pattern may lead to transient pathological behaviors.
-Maximizing convergence between these control mechanisms requires increasing the likelihood of both an intersection and a continuous path in conductance space.
-We hypothesize that this can be achieved by maximizing neuronal degeneracy (Whitacre,, 2010; Rathour and Narayanan,, 2019; Whitacre and Bender,, 2010; Prinz,, 2017; Goaillard and Marder,, 2021).
+As shown in Fig 4, it corresponds to the intersection of the strong bursting isocline (the target of controlled neuromodulation) and the homeostatic calcium setpoint.
+When no such intersection exists – e.g., under sodium channel blockade (Fig 5) – neither neuromodulated activity nor calcium homeostasis can be sustained.
+This may lead to pathological compensations, where one controller prioritizes calcium regulation at the expense of the target firing pattern [48, 16, 1].
+Additionally, discontinuities in neuromodulated firing patterns may induce transient pathological behaviors.
+Maximizing convergence between these control mechanisms requires increasing the probability of both an intersection and a continuous path in conductance space.
+We hypothesize that this can be achieved by maximizing neuronal degeneracy [54, 44, 53, 42, 15].
 Higher degeneracy increases the likelihood that the functional and dynamical role of one ion channel can be compensated by others, thereby preserving both the path and the intersection.
-This underscores the essential role of degeneracy in enabling robust neural functions that can be modulated.
-Finally, this study underscores the potential risks associated with pharmacological ion channel blockade (X Cubeddu,, 2016; DeWitt and Waksman,, 2004; Imbrici et al.,, 2018).
-If neuronal degeneracy is insufficient, such interventions may disrupt neuron ability to maintain robust neuromodulation.
-Instead, we propose targeting elements within the neuromodulation cascade—such as second messengers or other components of the control loop—as more reliable pharmacological strategies.
-Such interventions would preserve the dynamics of controlled neuromodulation, providing a more robust alternative to direct ion channel blockade.
+This highlights the crucial role of degeneracy in enabling robust, neuromodulated neural function.
+Finally, these findings underscore the potential risks associated with pharmacological ion channel blockade [5, 9, 21].
+If neuronal degeneracy is insufficient, such interventions may compromise the ability to sustain robust neuromodulation.
+As an alternative, we propose targeting elements within the neuromodulation cascade – such as second messengers or other components of the control loop – as more reliable pharmacological strategies.
+By preserving the dynamics of controlled neuromodulation, these interventions could offer a more robust alternative to direct ion channel blockade.
 Materials and methods
 Programming language
-The Julia programming language was used in this work Bezanson et al., (2017). Numerical integration was realized using DifferentialEquations.jl.
-Conductance-based models
+The Julia programming language was used in this work [3]. Numerical integration was realized using DifferentialEquations.jl.
+Conductance-based model
 For all experiments, single-compartment conductance-based models were employed.
-These models articulate an ordinary differential equation for the membrane voltage $V$, where $N$ ion channels are characterized as nonlinear dynamic conductances, and the phospholipid bilayer is represented as a passive resistor-capacitance circuit.
+These models articulate an ordinary differential equation for the membrane voltage $V$, where $N$ ion channels are characterized as nonlinear dynamic conductances, and the phospholipid bilayer is represented as a passive resistor-capacitor circuit.
 Mathematically, the voltage-current relationship of any conductance-based neuron model is expressed as follows:
@@
 Each ion channel conductance is nonlinear and dynamic, represented by $g_{\mathrm{ion}}(V,t)=\bar{g}_{\mathrm{ion}}m_{\mathrm{ion}}^{a}(V,t)h_{\mathrm{ion}}^{b}(V,t)$, where $m_{\mathrm{ion}}$ and $h_{\mathrm{ion}}$ are variables gated between 0 and 1, modeling the opening and closing gates of ion channels, respectively.
-Throughout this study, both the isolated crab STG neuron model Liu et al., (1998) and the adapted DA neuron model Qian et al., (2014) (where SK channels had been blocked to enable bursting) were employed.
+Throughout this study, the isolated crab STG neuron model of [25] was employed.
 The STG model consists of seven ion channels that operate on various time scales: fast sodium channels ($\bar{g}_{\mathrm{Na}}$); delayed-rectifier potassium channels ($\bar{g}_{\mathrm{Kd}}$); T-type calcium channels ($\bar{g}_{\mathrm{CaT}}$); A-type potassium channels ($\bar{g}_{\mathrm{A}}$); slow calcium channels ($\bar{g}_{\mathrm{CaS}}$); calcium controlled potassium channels ($\bar{g}_{\mathrm{KCa}}$); and H channels ($\bar{g}_{\mathrm{H}}$).
-The DA model consists of six ion channels that operate on various time scales: fast sodium channels ($\bar{g}_{\mathrm{Na}}$); delayed-rectifier potassium channels ($\bar{g}_{\mathrm{Kd}}$); L-type calcium channels ($\bar{g}_{\mathrm{CaL}}$); N-type calcium channels ($\bar{g}_{\mathrm{CaN}}$); ERG channels ($\bar{g}_{\mathrm{ERG}}$); and NMDA channels ($\bar{g}_{\mathrm{NMDA}}$).
-Note that, owing to the multicellular nature of NMDA channels, they were excluded from this study but were still used for simulations with baseline values.
 The homeostatic controller
-The homeostatic controller adjusts all conductances using an integration rule described in O’Leary et al., (2014).
+The homeostatic controller adjusts all conductances using an integration rule described in [38].
 Let $\bar{g}_{i}$ denote the conductance of ion channel $i\in[1,N]$, and $m_{i}$ the corresponding mRNA concentration.
@@
 $\tau_{g}\dot{\bar{g}}_{i}$ $=m_{i}-\bar{g}_{i},$
-where $[Ca^{+2}]$ and $[Ca^{+2}]_{\mathrm{target}}$ represent the intracellular calcium concentration and its target value, respectively.
-In essence, the calcium concentration error is integrated into the mRNA levels, which subsequently modulate the corresponding channel conductances.
-This regulation aims to bring the calcium concentration to its target value, as higher channel conductances (particularly calcium-conducting ones) facilitate greater calcium influx.
+where $Ca^{+2}]$ and $Ca^{+2}]_{\mathrm{target}}$ represent the calcium level and its target value, respectively.
+In essence, the calcium level error is integrated into the mRNA levels, which subsequently modulate the corresponding channel conductances.
+This regulation aims to bring the calcium level to its target value, as higher channel conductances (particularly calcium-conducting ones) facilitate greater calcium influx.
 Starting from any initial condition, this controller drives the cell along a line in conductance space known as the homogeneous scaling line.
@@
 This invariance ensures that the relative scaling of conductances is preserved during homeostatic compensation.
-Consequently, the cell’s conductance properties develop along the homogeneous scaling beam defined by these ratios.
+Consequently, the cell conductance properties develop along the homogeneous scaling beam defined by these ratios.
 The neuromodulation controller
-The neuromodulation controller adjusts a subset of $n$ conductances, denoted as $\bar{g}_{\mathrm{mod}}\in\mathbb{R}^{n}$ with $n<N$, using the algorithm introduced in Fyon et al., (2023).
+The neuromodulation controller adjusts a subset of $n$ conductances, denoted as $\bar{g}_{\mathrm{mod}}\in\mathbb{R}^{n}$ with $n<N$, using the algorithm introduced in [14].
 This controller employs the concept of Dynamic Input Conductances (DICs) to determine target values for $\bar{g}_{\mathrm{mod}}$ based on the input neuromodulator concentration, represented as $\bar{g}_{0}\left([\mathrm{nmod}]\right)\in\mathbb{R}^{n}$.
@@
 where $K_{p}$ and $K_{i}$ are the proportional and integral gains of the PI controller, respectively.
-In summary, the controller adjusts the modulated conductances $\bar{g}_{\mathrm{mod}}$ in response to the local concentration of neuromodulator, producing new reference values for these conductances to fine-tune the cell’s firing pattern—effectively neuromodulating it.
+In summary, the controller adjusts the modulated conductances $\bar{g}_{\mathrm{mod}}$ in response to the local concentration of neuromodulator, producing new reference values for these conductances to fine-tune the cell firing pattern – effectively neuromodulating it.
 This neuromodulator controller produces changes in conductance ratios that are normally preserved within the homeostatic controller.
@@
 DICs consist of three voltage-dependent conductances that separate according to timescales: one fast, one slow, and one ultraslow, denoted as $g_{\mathrm{f}}(V)$, $g_{\mathrm{s}}(V)$, and $g_{\mathrm{u}}(V)$, which can be computed as linear functions of the maximal conductance vector $\bar{g}_{\mathrm{ion}}\in\mathbb{R}^{N}$ of an $N$-channel conductance-based model at each voltage level $V$:
-$[g_{\mathrm{f}}(V);g_{\mathrm{s}}(V);g_{\mathrm{u}}(V)]=f_{\mathrm{DIC}}(V)=S(V)\cdot\bar{g}_{\mathrm{ion}}\,,$
-where $S(V)\in\mathbb{R}^{3\times N}$ is a sensitivity matrix that can be built by: $S_{i\,j}(V)=-\left(w_{i\,j}\cdot\frac{\partial\dot{V}}{\partial X_{j}}\frac{\partial X_{j,\infty}}{\partial V}\right)/g_{\mathrm{leak}}$, where $i$ denotes the timescale, $X_{j}$ are gating variables of the j-th channel of the considered model and $w_{i\,j}$ is a timescale-dependent weight which is computed as the logarithmic distance of the time constant of $X_{j}$ and the timescale $i$ (Drion et al., 2015a, ).
-While the complete curve of the DICs may be of interest, only its value at the threshold voltage $V_{\mathrm{th}}$ is used, as the values and signs of the DICs at $V_{\mathrm{th}}$ reliably determine the firing pattern Drion et al., 2015a .
+$g_{\mathrm{f}}(V);g_{\mathrm{s}}(V);g_{\mathrm{u}}(V)]=f_{\mathrm{DIC}}(V)=S(V)\cdot\bar{g}_{\mathrm{ion}}\,$,
+where $S(V)\in\mathbb{R}^{3\times N}$ is a sensitivity matrix that can be built by: $S_{i\,j}(V)=-\left(w_{i\,j}\cdot\frac{\partial\dot{V}}{\partial X_{j}}\frac{\partial X_{j,\infty}}{\partial V}\right)/g_{\mathrm{leak}}$, where $i$ denotes the timescale, $X_{j}$ are gating variables of the j-th channel of the considered model and $w_{i\,j}$ is a timescale-dependent weight which is computed as the logarithmic distance of the time constant of $X_{j}$ and the timescale $i$ [10].
+While the complete curve of the DICs may be of interest, only its value at the threshold voltage $V_{\mathrm{th}}$ is used, as the values and signs of the DICs at $V_{\mathrm{th}}$ reliably determine the firing pattern [10, 4].
 Thus, the following linear system $f_{\mathrm{DIC}}(V_{\mathrm{th}})=S(V_{\mathrm{th}})\cdot\bar{g}_{\mathrm{ion}}$ makes the link between ion channel conductances and neuronal activity.
-COMPETING INTEREST
-The authors declare no competing interests.
-FUNDING
-Arthur Fyon is a Research Fellow of the “Fonds de la Recherche Scientifique - FNRS”, supported by grant ASP-REN40024838.
-This work was supported by the Belgian Government through the Federal Public Service (FPS) Policy and Support (BOSA) grant NEMODEI (NEuroMOrphic Design of Embodied Intelligence).
-AUTHOR CONTRIBUTIONS STATEMENT
-A.F.
-and G.D.
-conceived and designed the research.
-A.F.
-and G.D.
-developed the model code.
-A.F.
-performed the simulations and analyzed the data.
-A.F.
-and G.D.
-interpreted the results.
-A.F.
-prepared figures and drafted the manuscript.
-A.F.
-and G.D.
-edited and revised the manuscript.
-A.F.
-and G.D.
-approved the final version of the manuscript.
-DATA AVAILABILITY
-All code and data can be found on the first author’s GitHub (https://github.com/arthur-fyon/BIOCONTROL_2025).
+Use of generative AI
+The ChatGPT-5 chatbot has been used to improve the syntax and grammar of several paragraphs in the manuscript.
+After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the content of the published article.
+Supporting information
+S1 Appendix
+Details and additional experiments We provide all the elements required to reproduce the results presented in this paper, as well as additional experiments and their results.
+This includes, for example, model equations and their parameter values.

### v3

@@
-Controlled neuromodulation naturally interacts with homeostasis to produce robust and modulable neuronal function
+Activity-dependent neuromodulation and calcium homeostasis cooperate to produce robust and modulable neuronal function
 Abstract
-Neurons rely on two interdependent mechanisms – homeostasis and neuromodulation – to maintain robust and adaptable functionality.
-Homeostasis stabilizes neuronal activity by adjusting ionic conductances, whereas neuromodulation dynamically modifies ionic properties in response to external signals carried by neuromodulators.
-Combining these mechanisms in conductance-based models often produces unreliable outcomes, particularly when sharp neuromodulation interferes with homeostatic tuning.
-This study explores how a biologically inspired neuromodulation controller can harmonize with homeostasis to ensure reliable neuronal function.
-Using computational models of stomatogastric ganglion and dopaminergic neurons, we demonstrate that controlled neuromodulation preserves neuronal firing patterns while homeostasis simultaneously maintains target intracellular calcium levels.
+Neurons rely on two interdependent mechanisms — homeostasis and neuromodulation — to maintain robust and adaptable functionality.
+Calcium homeostasis stabilizes neuronal activity by adjusting ionic conductances, whereas neuromodulation dynamically modifies ionic properties in response to external signals carried by neuromodulators.
+Combining these mechanisms in conductance-based models often produces unreliable outcomes, particularly when sharp neuromodulation interferes with calcium-homeostatic tuning.
+This study explores how a biologically inspired neuromodulation controller can harmonize with calcium homeostasis to ensure reliable neuronal function.
+Using computational models of stomatogastric ganglion and dopaminergic neurons, we demonstrate that controlled neuromodulation preserves neuronal firing patterns while calcium homeostasis simultaneously maintains target intracellular calcium levels.
 Unlike sharp neuromodulation, the neuromodulation controller integrates activity-dependent feedback through mechanisms mimicking G-protein-coupled receptor cascades.
@@
 Maximizing neuronal degeneracy enhances the likelihood of such intersections, enabling robust modulation and compensation for channel blockades.
-We further show that this controller pairing extends to network-level activity, reliably modulating central pattern generators rhythmic activity.
-This study highlights the complementary roles of homeostasis and neuromodulation, proposing a unified control framework for maintaining robust and adaptive neural activity under physiological and pathological conditions.
+We further show that this controller pairing extends to network-level activity, reliably modulating the rhythmic activity of central pattern generators.
+This study highlights the complementary roles of calcium homeostasis and neuromodulation, proposing a unified control framework for maintaining robust and adaptive neural activity under physiological and pathological conditions.
 Author summary
-Neurons must maintain stable activity while adapting to changing demands, relying on two mechanisms: homeostasis, which stabilizes activity, and neuromodulation, which adjusts behavior to external signals.
+Neurons must maintain stable activity while adapting to changing demands, relying on two mechanisms: calcium homeostasis, which stabilizes activity, and neuromodulation, which adjusts behavior to external signals.
 These mechanisms often interact, but their improper coordination can lead to dysfunction.
-This study uses computational models to show how controlled neuromodulation – mimicking biological feedback – can be harmonized with controlled homeostasis to ensure robust neuronal function.
-The system compensates for channel blockades and maintains critical activity patterns for both single neurons and networks even when faced with neuronal parameter variability.
+This study uses computational models to show how controlled neuromodulation — mimicking biological feedback — can be harmonized with calcium homeostasis to ensure robust neuronal function.
+The system compensates for channel blockades and maintains critical activity patterns for both single neurons and networks despite neuronal parameter variability.
 These insights not only deepen our understanding of neural robustness, but also suggest safer pharmacological strategies targeting neuromodulatory pathways, offering potential breakthroughs in treating neurological disorders without disrupting essential cellular mechanisms.
@@
 Among homeostatic mechanisms, intrinsic homeostatic plasticity adjusts ion channel conductances in response to intracellular calcium levels [43, 49].
-However, since both neuromodulation and homeostasis act on the same targets (ion channels) and may have competing objectives, understanding their interaction is essential [30, 31].
-We propose to study the interaction between neuromodulation and homeostasis using conductance-based models.
+However, since both neuromodulation and calcium homeostasis act on the same targets (ion channels) and may have competing objectives, understanding their interaction is essential [37, 38].
+We propose to study the interaction between neuromodulation and calcium homeostasis using conductance-based models.
 These models provide a biophysically accurate representation of neuronal dynamics by describing the electrical properties of the neuronal membrane.
 The membrane is modeled as an RC circuit, while ion channels are represented as nonlinear conductances that evolve according to experimentally derived activation and inactivation kinetics [28].
-Conductance-based models remain the most faithful mathematical representations of biological neurons, as they allow for the direct incorporation of ion channel dynamics extracted from experimental data, following the methodology first introduced by Hodgkin and Huxley [19].
+Conductance-based models remain the most faithful mathematical representations of biological neurons, as they allow direct incorporation of ion channel dynamics extracted from experimental data, following the methodology first introduced by Hodgkin and Huxley [24].
 Over the years, numerous conductance-based models have been developed to describe a variety of neuronal cell types across different species and brain regions [31, 52, 47, 64, 13, 29, 41].
@@
 STG neurons are key components of rhythmic circuits that generate motor patterns driving stomach muscle contractions in crustaceans.
-This model was chosen for its ability to reproduce diverse firing patterns – ranging from tonic spiking to bursting – and its rich set of ion channels, which enable degeneracy.
+This model was chosen for its ability to reproduce diverse firing patterns — ranging from tonic spiking to bursting — and its rich set of ion channels, which enable degeneracy.
 Specifically, the model includes classical fast sodium and delayed rectifier potassium channels, A-type and calcium-activated potassium channels, T-type and slow calcium channels, as well as hyperpolarization-activated H-channels.
-We model intrinsic homeostasis by using the homeostatic controller proposed by [38], which adjusts all ionic conductances up or down based on the neuron average intracellular calcium level.
-If the calcium level is lower than a target reference value, the controller increases ionic conductances according to homeostatic tuning rules, and conversely, decreases them if the calcium level is above the target.
+We model intrinsic homeostasis using the calcium-homeostatic controller proposed by [45], which adjusts all ionic conductances based on the deviation of the neuron average intracellular calcium level from a target set-point.
+We refer to this mechanism as calcium homeostasis throughout this work, to distinguish it from other forms of homeostatic plasticity.
+Specifically, if the calcium level is lower than a target reference value, the controller increases ionic conductances according to calcium-homeostatic tuning rules, and conversely, decreases them if the calcium level is above the target.
 Importantly, these rules ensure that conductances are tuned along a single axis, preserving their relative ratios and maintaining correlations between them.
@@
 Neuromodulation is modeled using two different approaches.
-On the one hand, we model neuromodulation as an acute change in target conductance values leading to a desired neuromodulated state in degenerate populations, which we call ”sharp neuromodulation”.
+On the one hand, we model neuromodulation as a hardwired change in target conductance values leading to a desired neuromodulated state in degenerate populations, which we call “sharp neuromodulation”.
 On the other hand, we implement the neuromodulation controller developed by [16], which is designed to enable robust modulation of neuronal activity.
-We call this approach ”controlled neuromodulation”.
+We call this approach “controlled neuromodulation”.
 One of its key properties is that it accounts for neuronal degeneracy, meaning that neurons with different ion channel compositions but similar activity patterns respond in a consistent manner to neuromodulation [17].
@@
 By leveraging degeneracy, the neuromodulation controller can apply the same control signal across a population of neurons with diverse conductance profiles, ensuring reliable and predictable neuromodulation.
-We first investigate the interaction between intrinsic homeostatic plasticity and neuromodulation at the single-neuron level and show that combining intrinsic homeostasis with either sharp or controlled neuromodulation leads to vastly different outcomes.
-Secondly, we provide a mechanistic explanation for these different outcomes.
+We first describe the calcium-homeostatic controller and the two neuromodulation approaches, sharp and controlled, highlighting their distinct dependence on neuronal activity.
+We then investigate their interaction at the single-neuron level and show that combining intrinsic homeostasis with either sharp or controlled neuromodulation leads to vastly different outcomes.
+Next, we provide a mechanistic explanation for these different outcomes.
 We then extend our analysis to rhythmic neuronal networks, where neuromodulation plays a critical role in adjusting oscillatory patterns.
 By carefully integrating homeostatic and neuromodulatory control mechanisms, we aim to develop a unified framework that combines the strengths of both approaches.
-This integrated controller enables robust modulation of neuronal activity while ensuring long-term stability, shedding light on how neuromodulation and homeostasis coexist in biological systems.
+This integrated controller enables robust modulation of neuronal activity while ensuring long-term stability, shedding light on how neuromodulation and calcium homeostasis coexist in biological systems.
 Results
-Interference between homeostasis and sharp neuromodulation leads to pathological outcomes
-We first examine the interplay between a state-of-the-art model of cellular homeostasis [38] and sharp neuromodulation [13].
-Here, sharp neuromodulation is defined as a direct, externally driven modification of a subset of ion channel maximal conductances, resulting in a change in neuronal spiking behavior.
+Calcium homeostasis and neuromodulation: two complementary regulatory mechanisms
+We begin by describing the two regulatory mechanisms that are central to this study: calcium homeostasis and neuromodulation.
+Both mechanisms act on ion channel conductances, but they differ fundamentally in their objectives, timescales, and dependence on neuronal activity.
+The calcium-homeostatic model [45] relies on intracellular calcium as a feedback signal, as illustrated in Fig 1A.
+The neuron continuously averages its intracellular calcium level over time and adjusts conductances accordingly: if calcium falls below a target level, ion channel transcription and translation increase, raising all conductance values, and vice versa.
+This phenomenon happens on a relatively slow timescale, which can last up to days [43].
+Importantly, the calcium-homeostatic tuning rules ensure that conductances are tuned along a single axis, preserving their relative ratios and maintaining correlations between them.
+Starting from any initial condition, this controller drives the cell along a line in conductance space known as the homogeneous scaling line.
+This self-tuning mechanism allows neurons to counteract perturbations while preserving stable activity.
+Fig 1: Schematic representations of calcium homeostasis, sharp and controlled neuromodulation.
+A.
+Schematic representation of calcium-homeostatic cascade.
+It senses intracellular calcium levels and regulates them by adjusting all conductances uniformly.
+B.
+Schematic representations of sharp (left) and controlled (right) neuromodulation.
+In sharp neuromodulation, neuromodulators bind to G-protein-coupled receptors, which directly affect subsets of ion channels independently of neuronal activity.
+In contrast, during controlled neuromodulation, G-protein-coupled receptors trigger complex signaling pathways that selectively affect subsets of ion channels in an activity-dependent way, hence producing a separate feedback loop over neuronal activity, acting in parallel with the calcium-homeostatic loop.
+Created in Bio Render.
+Fyon, A.
+(2026) https://BioRender.com/6q3irjy.
+In contrast, neuromodulation modifies a subset of ion channel conductances to reshape firing patterns.
+We consider two distinct implementations: sharp neuromodulation and controlled neuromodulation.
+Sharp neuromodulation is defined as a direct, hardwired, externally driven modification of a subset of ion channel maximal conductances, applied independently of neuronal activity (Fig 1B left).
+In this approach, neuromodulators bind to G-protein-coupled receptors, which directly affect subsets of ion channels without incorporating feedback from the neuron current activity state.
+Sharp neuromodulation, implemented using the method from [15], thus upsets the homeostasis setpoint, defined by the intracellular calcium target level.
+As calcium homeostasis subsequently acts to recover its calcium set point, there are no guarantees that the neuromodulated function is maintained as ion channel conductances are modified.
+Alternatively, we can take into account that, in biological systems, neuromodulation occurs through an intracellular cascade that indirectly links changes in neuromodulator concentration and changes in targeted conductance values using intermediate signals such as, e.g., cyclic adenosine monophosphate (cAMP) or intracellular calcium.
+We have recently shown that this indirect action of neuromodulation on conductance values could be viewed as an intracellular controller (Fig 1B right), which greatly improves the reliability of neuromodulatory action on highly degenerate populations [16].
+We call this mechanism “controlled neuromodulation” to contrast with the sharp neuromodulation described above.
+Unlike sharp neuromodulation, which imposes activity-independent conductance changes, the neuromodulation controller provides continuous and adaptive modulation in an activity-dependent manner, ensuring stable functional transitions.
+This approach is biologically inspired, as there is experimental evidence that neuromodulation acts in an activity-dependent manner [30, 54, 19, 18, 56, 37].
+Specifically, the neuromodulation controller models the role of G-protein-coupled receptors and second messengers in a non-specific way, embedding feedback within the metabolic neuromodulation cascade [33, 55].
+During controlled neuromodulation, G-protein-coupled receptors trigger complex signaling pathways that selectively affect subsets of ion channels in an activity-dependent way, hence producing a separate feedback loop over neuronal activity, acting in parallel with the calcium-homeostatic loop.
+At its core, the controller exploits Dynamic Input Conductances (DICs), a low-dimensional representation that links ion channel conductances to firing patterns (see Materials and Methods for details).
+DICs reduce the high-dimensional space of ion channel conductances to three voltage-dependent feedback gains, each associated with a distinct timescale (fast, slow, and ultraslow), whose values at threshold voltage reliably determine the firing pattern [11].
+The controller operates in two stages.
+First, a feedforward step uses this linear mapping to compute neuron-specific reference values for the modulated conductances, given the desired firing pattern and the current state of all other (unmodulated) conductances.
+Second, a feedback step in the form of a Proportional-Integral (PI) controller tracks these references by regulating ion channel expression.
+Because the feedforward step depends on the full conductance state of the neuron, the controller naturally adapts its output to each individual neuron within a degenerate population, producing different conductance adjustments that all converge to the same target activity.
+Importantly, the controller is agnostic to the specific neuromodulator identity and intracellular signaling cascade involved.
+This abstraction is justified by experimental evidence showing that diverse signaling pathways converge onto the same functional targets [37].
+For instance, in the STG, multiple neuromodulators acting through distinct G-protein-coupled receptors and second messenger pathways (involving cAMP or calcium-dependent cascades) all converge onto the same modulator-activated inward current $I_{\mathrm{MI}}$ [60].
+In this context, the feedforward step captures the dependence of neuromodulatory effects on the current activity state of the neuron [30], while the feedback step (PI controller) models the temporal integration inherent to intracellular signaling cascades, where second messenger concentrations gradually adjust in response to sustained receptor activation.
+Rather than modeling any particular cascade, our controller captures the functional principle common to these diverse pathways: an activity-dependent feedback loop that senses neuronal output and adjusts a targeted subset of conductances to reach a desired activity pattern.
+Sharp neuromodulation interferes with calcium homeostasis whereas controlled neuromodulation naturally leads to robust and modulable neuronal function
+Having established the two neuromodulation approaches, we now examine their interaction with calcium homeostasis.
 We specifically focus on a robust transition from tonic spiking to bursting in degenerate neuronal populations, i.e.
 populations composed of neurons that exhibit similar firing behaviors despite differences in underlying conductance profiles.
-To explore this interaction, we applied sharp neuromodulation to a population of degenerate neurons generated using the dataset from [13] (Fig 1).
-Sharp neuromodulation consisted of directly modifying the maximal conductances of A-type potassium ($\bar{g}_{\mathrm{A}}$) and slow calcium ($\bar{g}_{\mathrm{CaS}}$) channels to transition neurons from tonic spiking to bursting.
-We then allowed homeostatic compensation to unfold.
-Since bursting increases overall neuronal activity, intracellular calcium levels rise, triggering a decrease in all conductances according to homeostatic tuning rules.
-Three representative conductance trajectories ($N=200$) are shown in Fig 1 (top), while Fig 1 (bottom) illustrates calcium dynamics across the population (note that to facilitate observation of both mechanisms within a single trace, the homeostatic controller was accelerated compared to its original implementation).
-Despite initially similar firing patterns and a similar outcome of sharp neuromodulation, homeostatic compensation produced highly variable and often unreliable outcomes in terms of modulated activity, while still regulating intracellular calcium.
-Depending on the initial conductance state, neuromodulation led to new states where homeostasis could either maintain functional activity, result in pathological behavior, or drive neurons into unphysiological regimes.
-This finding underscores a fundamental property of degeneracy: neurons with different underlying conductances can exhibit similar activity yet respond differently to perturbations, such as neuromodulation.
-Fig 1: Interference between homeostasis and sharp neuromodulation leads to pathological outcomes.
-Time evolution of all conductances in the STG model, displayed on a logarithmic scale, during homeostasis with sharp neuromodulation (instantaneous changes in $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$ starting at the dashed blue line) for a degenerate population of 200 models (top).
-The corresponding mean intracellular calcium concentration over time (bottom) shows the target value (red dash-dotted line) is maintained both before and after sharp neuromodulation.
-To understand the source of this unreliability, it is necessary to examine both mechanisms in more detail.
-The homeostatic model [38] relies on intracellular calcium as a feedback signal.
-The neuron continuously averages its intracellular calcium level over time and adjusts conductances accordingly: if calcium falls below a target level (red dash-dotted line in Fig 1 bottom), ion channel transcription and translation increase, raising all conductance values, and vice versa.
-This phenomenon happens on a relatively slow timescale, which can last up to days [36].
-In contrast, sharp neuromodulation, implemented using the method from [13], directly modifies a subset of ion channel conductances, which upsets the homeostasis setpoint, defined by the intracellular calcium target level.
-As homeostasis kicks in to recover its calcium set point, there are no guarantees that the neuromodulated function is maintained as ion channel conductances are modified, which often leads to pathological or unphysiological outcomes.
-This type of interference was anticipated in [38], where the authors demonstrated that homeostatic regulation can produce diverse responses to perturbations such as ion channel deletions.
+We applied both types of neuromodulation to a population of degenerate neurons generated using the dataset from [15], modifying the maximal conductances of A-type potassium ($\bar{g}_{\mathrm{A}}$) and slow calcium ($\bar{g}_{\mathrm{CaS}}$) channels to transition neurons from tonic spiking to bursting, and then allowed calcium-homeostatic compensation to unfold.
+In all illustrative simulations (Fig 2), the time constant of the calcium-homeostatic controller was reduced so that the dynamics of both neuromodulation and calcium-homeostatic compensation could be observed within a single trace, without affecting the qualitative behavior of the system.
+Fig 2: Sharp neuromodulation interferes with homeostasis whereas controlled neuromodulation leads to robust and modulable neuronal function.
+A.
+Time evolution of all conductances in the STG model, displayed on a logarithmic scale, during calcium homeostasis with sharp neuromodulation (instantaneous changes in $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$ starting at the dashed blue line) for a degenerate population of $N=200$ neuron models.
+B.
+The corresponding mean intracellular calcium concentration over time shows the target value (red dash-dotted line) is maintained both before and after sharp neuromodulation.
+C.
+Three representative traces of neurons within the population are highlighted (top) before neuromodulation ($t_{\text{before}}$), right after ($t_{\text{after}}$) and after calcium-homeostatic compensation ($t_{\text{compensated}}$); along with the activity distributions of the population at these three time points (bottom).
+On the one hand, before neuromodulation, the population exhibits tonic firing, and hence the firing frequency (defined as the reciprocal of the interspike interval) is used as the activity metric.
+On the other hand, after neuromodulation, the population exhibits bursting, and hence the burstiness is used as the activity metric.
+Burstiness is defined as in [14], i.e., the product of the intraburst frequency (reciprocal of the interspike interval within a burst), the interburst frequency (reciprocal of the interburst interval), and the number of spikes per burst.
+A burstiness of zero means that the neuron is not producing multi-spike bursts (silence, tonic firing, or single-spike bursters).
+D.
+Same as panel A, but with controlled neuromodulation (controlled changes in $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$ starting at the dashed blue line).
+E.
+Same as panel B, but with controlled neuromodulation.
+F.
+Same as panel C, but with controlled neuromodulation.
+We first test the interaction of calcium homeostasis and sharp neuromodulation (Fig 2A-C).
+All conductance trajectories ($N=200$) are shown in Fig 2A, while Fig 2B illustrates calcium dynamics across the population.
+Specific time evolution curves for single neurons within the population can be found in Supporting Information, S1 Appendix.
+Despite initially similar firing patterns ($t_{\text{before}}$) and a similar outcome of sharp neuromodulation ($t_{\text{after}}$), calcium-homeostatic compensation produced highly variable and often unreliable outcomes in terms of modulated activity ($t_{\text{compensated}}$), while still regulating intracellular calcium (Fig 2C).
+Depending on the initial conductance state, neuromodulation led to new states where calcium homeostasis could either maintain functional activity, result in pathological behavior (loss of bursting, i.e., a burstiness of 0), or drive neurons into unphysiological regimes.
+This finding underscores a fundamental property of degeneracy: neurons with different underlying conductances can exhibit similar activity yet respond differently to perturbations, such as sharp neuromodulation.
+Note that these differences in channel conductances are consistent with biological measurements: different channel types have effective conductances spanning several orders of magnitude [50, 45].
+Moreover, the effective conductance of a given ion channel type can vary severalfold within a single neuron type [57].
+The variability in the dataset from [15] is inspired by these biological measurements and slightly exceeds the observed range to test the robustness of our analysis.
+This type of interference was anticipated in [45], where the authors demonstrated that calcium-homeostatic regulation can produce diverse responses to perturbations such as ion channel deletions.
 Depending on the initial conditions, a deletion may either preserve function, disrupt it, or lead to pathological compensation.
-Our results extend this idea by showing that even physiological disturbances, such as neuromodulation, can induce similar unpredictability when coupled with homeostatic regulation.
-Controlled neuromodulation and homeostasis naturally lead to robust and modulable neuronal function
-In biological systems, neuromodulation occurs through an intracellular cascade that indirectly links changes in neuromodulation concentration and changes in targeted conductance values using an intermediate signal such as e.g.
-cyclic amp.
-We have recently shown that this indirect action of neuromodulation on conductance values could be viewed as an intracellular controller (Fig 2A, top), which largely improves reliability of neuromodulatory action on highly degenerate populations [14].
-We call this mechanism ”controlled neuromodulation” to contrast with the sharp neuromodulation described above.
-Unlike sharp neuromodulation, which imposes instantaneous conductance changes, the neuromodulation controller provides continuous and adaptive modulation, ensuring stable functional transitions.
-Fig 2: Controlled neuromodulation and homeostasis naturally lead to robust and modulable neuronal function.
-A.
-Schematic representation of neuromodulation (left) and homeostasis (right) cascades.
-Neuromodulators bind to G-protein receptors, triggering complex signaling pathways that selectively affect subsets of ion channels.
-In contrast, homeostasis senses intracellular calcium levels and regulates them by adjusting all conductances uniformly.
-B.
-Time evolution of all conductances in the STG model, displayed on a logarithmic scale, during homeostasis with controlled neuromodulation (controlled changes in $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$) for a degenerate population of 200 models (left).
-The corresponding mean intracellular calcium concentration over time (right) shows the target value (red dash-dotted line) is maintained both before and after controlled neuromodulation.
-Using the same protocol as for sharp neuromodulation, we now study the interaction between controlled neuromodulation and homeostasis.
-Fig 2B illustrates the transition from tonic spiking to bursting in a degenerate population of STG neuron models, identical to those in Fig 1, but when using controlled neuromodulation instead of sharp neuromodulation.
-Initially, all neurons exhibit tonic firing, with homeostasis tuning conductances along a tonic spiking homogeneous scaling line to maintain a target calcium level.
-Midway through the simulation, controlled neuromodulation is applied, gradually adjusting $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$ to induce bursting.
+Our results extend this idea by showing that even physiological disturbances, such as neuromodulation, can induce similar unpredictability when coupled with calcium-homeostatic regulation.
+Using controlled neuromodulation on the same population dramatically changes the outcome (Fig 2D-F).
+Fig 2D-E illustrates the transition from tonic spiking to bursting in the same degenerate population of STG neuron models as in Fig 2A-C.
+Specific time evolution curves for single neurons within the population can be found in Supporting Information, S1 Appendix.
+Initially, all neurons exhibit tonic firing, with calcium homeostasis tuning conductances along a tonic spiking homogeneous scaling line to maintain a target calcium level ($t_{\text{before}}$ in Fig 2F).
+Midway through the simulation, controlled neuromodulation is applied, controlling $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$ to induce bursting ($t_{\text{after}}$ in Fig 2F).
 As in the case of sharp neuromodulation, this transition increases intracellular calcium levels, leading to an adaptive shift in conductance values.
-Homeostatic regulation then readjusts overall conductance magnitudes, but here the interaction between homeostasis and controller neuromodulation ensures that the newly established neuromodulated state is maintained.
-Ultimately, the population reaches a stable configuration where both the neuromodulated function and intracellular calcium homeostasis are fulfilled.
-To demonstrate the generality of this approach, we applied the same framework to a midbrain dopaminergic neuron model adapted from [43] (see Supporting Information, S1 Appendix).
-As in the STG model, a transition from tonic spiking (pacemaking) to strong bursting was induced by modulating N-type and L-type calcium conductances ($\bar{g}_{\mathrm{CaN}}$ and $\bar{g}_{\mathrm{CaL}}$) through controlled neuromodulation.
-Homeostatic compensation continuously regulated calcium levels, ensuring a reliable and robust outcome despite initial conductance degeneracy within the neuronal population.
-Controlled neuromodulation tunes the direction of homeostasis-driven changes in conductance values to ensure stable neuromodulated function during homeostatic regulation
-To study the mechanisms underlying the constructive interaction between controlled neuromodulation and homeostasis, we plot the conductance trajectories of the simulations from Fig 1 in the plane of modulated conductances, $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$, during tonic spiking homeostatic compensation for both sharp and controlled neuromodulation (Fig 3A).
-Initially, as calcium levels remain below the threshold, homeostasis increases the modulated conductances along a homogeneous scaling direction characteristic of tonic spiking.
-This direction corresponds to a line intersecting the origin, with variability due to conductance degeneracy within the population, and its slope is determined by the relative time constants of mRNA transcription in the homeostasis model, with steady-state analysis predicting $\bar{g}_{i}/\bar{g}_{j}\approx\tau_{m_{j}}/\tau_{m_{i}}$ [38].
+Calcium-homeostatic regulation then readjusts overall conductance magnitudes, but here the interaction between calcium homeostasis and controlled neuromodulation ensures that the newly established neuromodulated state is maintained.
+Ultimately, the population reaches a stable configuration where both the neuromodulated function and intracellular calcium homeostasis are fulfilled ($t_{\text{compensated}}$ in Fig 2F).
+The differences between sharp and controlled neuromodulation can also be highlighted through neuromodulator washout, i.e., removal of the neuromodulatory input.
+At both $t_{\text{after}}$ and $t_{\text{compensated}}$, washout recovers tonic spiking across the population under both sharp and controlled neuromodulation, though for fundamentally different reasons and with different levels of fidelity.
+Under sharp neuromodulation, calcium-homeostatic compensation continuously drives conductances along the native tonic spiking scaling direction, so that tonic spiking is the only remaining stable activity pattern.
+Recovery of spiking upon washout is therefore expected, as the system has no other stable attractor.
+However, when washout occurs at $t_{\text{compensated}}$, the conductance distribution of the population does not exactly match its pre-neuromodulation state: because calcium homeostasis has been compensating for the sharp neuromodulatory perturbation throughout, the resulting conductance profiles are subtly altered, leading to a slightly different distribution of firing frequencies across the population.
+Although the qualitative behavior (tonic spiking) is preserved, this quantitative drift implies that repeated cycles of sharp neuromodulation and washout could progressively erode the fidelity of neuronal function, as each cycle introduces small but cumulative changes in the conductance landscape.
+In contrast, under controlled neuromodulation, the system can maintain any target activity through activity-dependent feedback.
+Upon washout, the neuromodulatory reference is removed, and the system returns to tonic spiking because the calcium-homeostatic controller regains sole authority over conductance tuning.
+Because the controlled neuromodulation mechanism continuously coordinates with calcium homeostasis rather than competing with it, the pre-neuromodulation conductance distribution is faithfully recovered upon washout.
+Controlled neuromodulation tunes the direction of calcium homeostasis-driven changes in conductance values to ensure stable neuromodulated function during calcium-homeostatic regulation
+To study the mechanisms underlying the constructive interaction between controlled neuromodulation and calcium homeostasis, we plot the conductance trajectories of the simulations from Fig 2 in the plane of modulated conductances, $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$, during tonic spiking calcium-homeostatic compensation for both sharp and controlled neuromodulation (Fig 3A).
+Initially, as calcium levels remain below the threshold, calcium homeostasis increases the modulated conductances along a homogeneous scaling direction characteristic of tonic spiking.
+This direction corresponds to a line intersecting the origin, with variability due to conductance degeneracy within the population, and its slope is determined by the relative time constants of mRNA transcription in the calcium homeostasis model, with steady-state analysis predicting $\bar{g}_{i}/\bar{g}_{j}\approx\tau_{m_{j}}/\tau_{m_{i}}$ [45].
 When neuromodulation is applied, all conductances are adjusted to reshape their ratios and induce bursting (Fig 3B).
@@
 This modulation leads to a sharp increase in intracellular calcium levels.
-Once neuromodulation ceases, the long-term outcome depends on the type of neuromodulation applied (Fig 3C).
-Fig 3: Controlled neuromodulation tunes the direction of homeostasis-driven changes in conductance values to ensure stable neuromodulated function during homeostatic regulation.
+After neuromodulation is applied, the long-term outcome depends on the type of neuromodulation used (Fig 3C-D).
+Fig 3: Controlled neuromodulation tunes the direction of homeostasis-driven changes to ensure stable neuromodulated function.
 A.
 Trajectories of the STG population from Fig 2 in the modulated conductance space during homogeneous scaling in tonic spiking only.
-As expected from homeostasis, all neurons move in the same direction.
+As expected from calcium homeostasis, all neurons move in the same direction.
 White circles represent the initial conditions at the start of the step, and the green triangle marks the final conditions.
-B. Same as panel A, but during the brief period when neuromodulation is active. All trajectories are parallel to one another.
+B.
+Same as panel A, but during the brief period when neuromodulation is active.
+All trajectories are parallel to one another.
 C.
-Same as panels A and B, but during the homeostasis phase following neuromodulation.
-The outcome depends on the type of neuromodulation: sharp (left) or controlled (right).
-Sharp neuromodulation results in non-robust outcomes, whereas controlled neuromodulation preserves robust neural function.
-Under sharp neuromodulation, homeostasis operates alone in the bursting phase, following the scaling direction characteristic of its reference, i.e.
-unmodulated, state set by the relative time constants of mRNA transcription in the homeostasis model (Fig 3C, left).
-This results in a decrease in calcium levels, but as shown in Fig 1, the response is unreliable and often undesirable.
-Specifically, for most neurons, the homeostatic controller reduces $\bar{g}_{\mathrm{A}}$ toward negative values (saturated at 0 to prevent model instability) to compensate for $\bar{g}_{\mathrm{CaS}}$, driving the neuron away from robust bursting and leading to heterogeneous activity changes across the population.
-In contrast, controlled neuromodulation allows the homeostatic and neuromodulation controllers to act together, maintaining the new conductance ratios associated with bursting (Fig 3C, right).
-As a result, homeostatic regulation follows a new homogeneous scaling trajectory aligned with bursting conductance ratios.
+Same as panels A and B, but during the calcium homeostasis compensation phase with sharp neuromodulation, resulting in non-robust outcomes.
+D.
+Same as panel C, but with controlled neuromodulation, resulting in function-preserving robust outcomes.
+Under sharp neuromodulation, calcium homeostasis operates alone in the bursting phase, following the scaling direction characteristic of its reference, i.e.
+unmodulated, state set by the relative time constants of mRNA transcription in the calcium homeostasis model (Fig 3C).
+This results in a decrease in calcium levels, but as shown in Fig 2A-C, the response is unreliable and often undesirable, resulting in a loss of robust bursting across the population (bursting is either lost entirely, sustained with substantially altered properties, or replaced by irregular activity).
+Specifically, for most neurons, the calcium-homeostatic controller reduces $\bar{g}_{\mathrm{A}}$ toward negative values (saturated at 0 to remain in physiological ranges and prevent model instability) to compensate for $\bar{g}_{\mathrm{CaS}}$, driving the neuron away from robust bursting and leading to heterogeneous activity changes across the population (violin plots in Fig 3C).
+In contrast, controlled neuromodulation allows the homeostatic and neuromodulation controllers to act together, maintaining the new conductance ratios associated with bursting (Fig 3D).
+As a result, calcium-homeostatic regulation follows a new homogeneous scaling trajectory aligned with bursting conductance ratios.
 These trajectories converge toward the origin, as in tonic spiking, but with a distinct direction determined by neuromodulation.
 Due to initial conductance degeneracy, the exact scaling varies across the population.
-The shift between the two homogeneous scalings – tonic spiking and bursting – emerges from neuromodulation, ensuring that the desired neuromodulated function is maintained.
-The incompatibility of sharp neuromodulation with homeostasis can be further analyzed using a schematic representation of the modulated conductance plane (Fig 4A).
+The shift between the two homogeneous scalings — tonic spiking and bursting — emerges from neuromodulation, ensuring that the desired neuromodulated function is maintained (violin plots in Fig 3D).
+The incompatibility of sharp neuromodulation with calcium homeostasis can be further analyzed using a schematic representation of the modulated conductance plane (Fig 4A).
 In this plane, different combinations of $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$ can yield similar calcium levels, forming calcium isoclines (green dashed lines).
@@
 While calcium isoclines are parallel, activity isoclines radiate from the origin and rotate around it.
-This organization reflects homeostatic tuning rules: different conductance ratios correspond to distinct firing modes [38].
-Initially, the neuron operates at the intersection of a target spiking isocline and a target calcium isocline.
-When sharp neuromodulation is applied, conductances shift to a new state corresponding to strong bursting, moving the neuron to a higher calcium isocline.
-Homeostasis then acts to restore calcium levels, but because it retains the ”native” tonic spiking scaling direction, it alters conductances along the spiking isocline rather than preserving the newly induced bursting state.
+This organization reflects calcium-homeostatic tuning rules: different conductance ratios correspond to distinct firing modes [45].
+Initially, the neuron operates at the intersection of a target spiking isocline and a target calcium isocline (at $t_{\text{before}}$).
+When sharp neuromodulation is applied (black arrow in Fig 4A), conductances shift to a new state corresponding to strong bursting, moving the neuron to a higher calcium isocline.
+Calcium homeostasis then acts to restore calcium levels (solid red arrow in Fig 4A), but because it retains the “native” tonic spiking scaling direction (dashed red arrow in Fig 4A), it alters conductances along the spiking isocline rather than preserving the newly induced bursting state.
 As a result, the neuron drifts away from the bursting isocline, potentially disrupting its function.
-This explains why sharp neuromodulation leads to unreliable outcomes when coupled with homeostasis.
+This explains why sharp neuromodulation leads to unreliable outcomes when coupled with calcium homeostasis.
 Fig 4: Schematic explanation of why controlled neuromodulation integrates better with homeostasis compared to sharp neuromodulation.
 A.
-Schematic of the modulated conductance space for homeostasis combined with sharp neuromodulation.
-Initially (white circles), the neuron spikes at the target calcium level, located at the intersection of the firing pattern isocline (dashed blue lines) and the calcium level isocline (dashed green lines).
-Following sharp neuromodulation to a new firing pattern (red diamond), calcium levels sharply increase.
-Homeostasis reduces calcium (solid red arrow) by moving along the same direction as before neuromodulation (dashed red arrow), causing the neuron to deviate from the neuromodulated firing pattern isocline.
+Schematic of the modulated conductance space for calcium homeostasis combined with sharp neuromodulation.
+Initially at $t_{\text{before}}$ (green triangles), the neuron spikes at the target calcium level, located at the intersection of the firing pattern isocline (dashed blue lines) and the calcium level isocline (dashed green lines).
+Following sharp neuromodulation (black arrow) to a new firing pattern at $t_{\text{after}}$ (blue diamond), calcium levels sharply increase.
+Homeostasis reduces calcium (solid red arrow) by moving along the same direction as before neuromodulation (dashed red arrow), causing the neuron to deviate from the neuromodulated firing pattern isocline at $t_{\text{compensated}}$ (red star).
 B.
 Same as panel A, but with controlled neuromodulation.
-The initial steps are identical, but after neuromodulation, as homeostasis acts to reduce calcium, controlled neuromodulation simultaneously adjusts to keep the neuron on the neuromodulated firing pattern isocline (sawtooth red arrow).
-Because controlled neuromodulation operates on a much faster timescale than homeostasis, the neuron remains on the neuromodulated firing pattern isocline throughout.
-To reconcile homeostasis and neuromodulation and achieve robust functional modulation, we leverage their distinct timescales.
-Neuromodulation controls specific ionic conductances, while homeostasis adjusts all conductances over a slower timescale to regulate calcium levels.
+The initial steps are identical, but after neuromodulation, as calcium homeostasis acts to reduce calcium, controlled neuromodulation simultaneously adjusts to keep the neuron on the neuromodulated firing pattern isocline (sawtooth red arrow), as controlled neuromodulation is activity-dependent.
+Because controlled neuromodulation operates on a much faster timescale than calcium homeostasis, the neuron remains on the neuromodulated firing pattern isocline throughout.
+Black arrow: trajectory of sharp (A) or controlled (B) neuromodulation.
+Solid red arrow: trajectory of calcium-homeostatic compensation.
+Dashed red arrow: direction of the calcium-homeostatic compensation in tonic firing (A).
+To reconcile calcium homeostasis and neuromodulation and achieve robust functional modulation, we leverage their distinct timescales.
+Neuromodulation controls specific ionic conductances, while calcium homeostasis adjusts all conductances over a slower timescale to regulate calcium levels.
 In the modulated conductance plane (Fig 4B), the neuron initially follows the same trajectory as in Fig 4A.
-However, under controlled neuromodulation, both controllers remain active.
-When homeostasis lowers calcium levels and shifts the neuron away from the bursting isocline, the neuromodulatory controller counteracts this deviation, bringing the neuron back toward the bursting isocline and preserving the desired function.
-Because the neuromodulatory controller operates orders of magnitude faster than the homeostatic controller, its rapid corrections render homeostatic perturbations negligible.
+However, under controlled neuromodulation, both controllers remain active in an activity-dependent way.
+When calcium homeostasis lowers calcium levels and shifts the neuron away from the bursting isocline, the neuromodulatory controller counteracts this deviation, bringing the neuron back toward the bursting isocline and preserving the desired function.
+Because the neuromodulatory controller operates orders of magnitude faster than the calcium-homeostatic controller, its rapid corrections render homeostatic perturbations negligible.
 This interaction constrains changes in conductance ratios along the target activity isocline (strong bursting in this case).
 Ultimately, the neuron stabilizes at the intersection of the target calcium and target activity isoclines, where both controllers achieve their objectives.
-Controlled neuromodulation and homeostasis ensure the preservation of function under physiologically recoverable disturbances
+The results above were obtained with controlled neuromodulation acting on A-type potassium ($\bar{g}_{\mathrm{A}}$) and slow calcium ($\bar{g}_{\mathrm{CaS}}$) conductances.
+However, because the neuromodulation controller is agnostic to the specific neuromodulator and signaling cascade, we also conducted experiments in which H-type channels ($\bar{g}_{\mathrm{H}}$) are modulated, given that these channels are well known to be sensitive to cAMP levels [1].
+An analogous figure to Fig 2, with neuromodulation acting on H-type channels, can be found in Supporting Information, S1 Appendix.
+These experiments yield similar results: failure under sharp neuromodulation and robust function under controlled neuromodulation.
+This confirms that the observations are not specific to the particular subset of modulated channels, but rather emerge from the activity-dependent control action itself.
+Controlled neuromodulation and calcium homeostasis ensure the preservation of function under physiologically recoverable disturbances
 By employing this tandem of controllers, neuronal function becomes robust to channel blockade, provided that the channel deletion is compensable.
-While homeostasis alone may lead to unreliable recovery, incorporating controlled neuromodulation stabilizes the neuronal response to blockade, even in highly degenerate neuronal populations.
-Furthermore, if the blocked channel function can be physiologically recovered – meaning other channels can substitute for its role – there exists an optimal combination of neuronal feedback gains that preserves the function.
-In essence, when a channel is blocked, a specific concentration of neuromodulators can maintain the function.
-For example, in Fig 5, various channel blockades are applied to a degenerate population of STG conductance-based models equipped with the combined homeostasis/neuromodulation tandem.
+While calcium homeostasis alone may lead to unreliable recovery, incorporating controlled neuromodulation stabilizes the neuronal response to blockade, even in highly degenerate neuronal populations.
+Furthermore, if the function of the blocked channel can be physiologically recovered, meaning that other channels can substitute for its role, there exists an optimal combination of neuronal feedback gains (i.e., controlled neuromodulation input) that preserves the function.
+In essence, when a channel is blocked and its function is compensable, a specific neuromodulator concentration (or cocktail of neuromodulators) can maintain the neuronal function.
+For example, in Fig 5, various channel blockades are applied to a degenerate population of STG conductance-based models equipped with the combined calcium homeostasis/neuromodulation tandem.
 The target behavior of the population is regular bursting.
 In Fig 5A, H-type channels are blocked.
-Here, the function remains intact even immediately after the blockade, without requiring compensation.
-The combined action of homeostasis and controlled neuromodulation then ensures that intracellular calcium levels are restored on a population-wide scale, maintaining the function.
+Here, the function remains intact even immediately after the blockade, without requiring compensation (violin plots in Fig 5A right).
+The combined action of calcium homeostasis and controlled neuromodulation then ensures that intracellular calcium levels are restored on a population-wide scale, maintaining the function.
 Fig 5: Controlled neuromodulation and homeostasis ensure the preservation of function under physiologically recoverable disturbances.
 A.
-Time evolution of all conductances in the STG model, displayed on a logarithmic scale, during homeostasis with controlled neuromodulation (controlled changes in $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$ starting at the dashed blue line) for a degenerate population of 200 models with H-type ion channel blockade (left).
+Time evolution of all conductances in the STG model, displayed on a logarithmic scale, during calcium homeostasis with controlled neuromodulation (controlled changes in $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$) for a degenerate population of $N=200$ neuron models with H-type ion channel blockade (left).
 The corresponding distribution of behaviors before the blockade, immediately after the blockade, and following compensation (right) demonstrates the effect of the blockade.
 In this case, no neurons lose bursting due to the channel blockade or compensation.
-B. Same as panel A, but with a CaT channel blockade. The blockade causes half of the neurons to lose bursting; however, all neurons recover bursting with compensation.
-C. Same as panel A, but with a KCa channel blockade. The blockade causes more than half of the neurons to lose bursting, but nearly all recover bursting with compensation.
+B.
+Same as panel A, but with a CaT channel blockade.
+The blockade causes half of the neurons to lose bursting; however, all neurons recover bursting with compensation.
+C.
+Same as panel A, but with a KCa channel blockade.
+The blockade causes more than half of the neurons to lose bursting, but nearly all recover bursting with compensation, although the burstiness of the population remains altered.
 D.
@@
 In Fig 5B and C, T-type calcium and calcium-activated potassium channels are blocked, respectively.
-In these cases, the blockade largely disrupts bursting.
-However, allowing the homeostasis/neuromodulation controllers to compensate restores function in most neurons, albeit with some displaying irregular bursting.
+In these cases, the blockade largely disrupts bursting (violin plots in Fig 5B-C right).
+However, allowing the calcium homeostasis/neuromodulation controllers to compensate restores function in most neurons, albeit with some displaying irregular bursting or reduced burstiness following compensation for the KCa channel blockade (violin plot in Fig 5C far right).
 Additionally, intracellular calcium levels stabilize at their target values.
-This recovery is possible because the actions of H-type, T-type calcium, and calcium-activated potassium channels are degenerate, allowing their function to be compensated by other channels in the cell.
+Despite operating at different timescales and membrane potentials, and thus producing different acute physiological effects upon blockade, recovery is possible for these channels because their functions overlap with those of other channels in the model, i.e., they exhibit degeneracy [17].
 The STG neuron model includes numerous channels operating on slow and ultraslow timescales, enabling compensation over these timescales.
-Notably, only the loss of the calcium-activated potassium channel has been compensated by adjusting neuronal feedback gains – equivalent to modifying neuromodulator concentrations – as in control conditions.
+The function of calcium-activated potassium channels is less degenerate: its role cannot be fully compensated, as reflected in the reduced population burstiness, although the bursting function itself is maintained.
+However, even when the blockade is theoretically compensable, the control input may need to be adjusted.
+In the case of the KCa channel blockade, the loss was compensated by tuning the neuronal feedback gains, which corresponds to applying a different neuromodulator concentration to partially recover the function.
+This highlights that, even when compensation is theoretically possible, the neuron may not reach the compensated state without adjusting the neuromodulator cocktail.
 In contrast, the blockade of fast transient sodium channels is irrecoverable (Fig 5D), as these channels are essential for spike initiation on a fast timescale.
@@
 CPGs are self-organized neuronal circuits composed primarily of bursting neurons connected through inhibitory synapses, capable of generating rhythmic output without external input [68, 21].
-A well-studied example is the circuit responsible for the pyloric and gastric mill rhythms in crustaceans, located in the stomatogastric ganglion.
-This network produces two distinct rhythms: the fast pyloric rhythm and the slow gastric mill rhythm.
+A well-studied example is the circuit responsible for the pyloric and gastric mill rhythms in crustaceans, located in the STG.
+This network produces two distinct rhythms: the fast pyloric rhythm and the neuromodulated slow gastric mill rhythm.
 These rhythmic outputs are essential for coordinating muscle contractions in the stomach for digestion, among other functions.
@@
 Half-center oscillators are formed by two bursting neurons connected through mutual inhibition, allowing them to fire anti-phase bursts.
-In this model, the purple and blue neurons constitute the fast half-center oscillator (representing the pyloric rhythm), while the green and yellow neurons form the slow half-center oscillator (representing the gastric mill rhythm).
+In this model, the purple and blue neurons constitute the fast half-center oscillator (representing the pyloric rhythm), while the green and yellow neurons form the neuromodulated slow half-center oscillator (representing the gastric mill rhythm).
 The red neuron acts as the central element, also in bursting mode.
@@
 A.
-Control condition of the simplified pyloric/gastric mill rhythm network (left), inspired from [18].
+Control condition of the simplified pyloric/gastric mill rhythm network composed of 5 neurons (N1-N5) (left), inspired by [22].
 Without neuromodulation, the fast rhythm is active (top two traces, middle), while the slow rhythm is inactive (bottom two traces, middle).
-In this case, all neurons in the network exhibit degenerate conductances (right).
 B.
+Initial conductance values for the five neurons of the network (N1-N5, color coded).
+Multiplier coefficients have been applied to all conductances to facilitate visualization in a non-logarithmic graph, allowing direct comparison of the different initial conductance values across all neurons, as in [50].
+C.
 Schematic representation of the activated slow rhythm network (left), modulated by either sharp neuromodulation (middle) or controlled neuromodulation (right).
-With sharp neuromodulation, the slow rhythm fails due to its unreliability, whereas controlled neuromodulation sustains the slow rhythm.
-In this study, the tandem of controllers reliably induces the gastric mill rhythm, whereas homeostasis combined with sharp neuromodulation does not.
+With sharp neuromodulation, the slow rhythm fails to be sustained due to its unreliability, whereas controlled neuromodulation sustains the slow rhythm.
+In this study, the tandem of controllers reliably induces the gastric mill rhythm, whereas calcium homeostasis combined with sharp neuromodulation does not.
 Under control conditions (Fig 6A), i.e., in the absence of neuromodulatory input to the gastric mill rhythm, only the pyloric rhythm remains active, with the central neuron bursting continuously.
-Notably, conductance variability is observed within the network (Fig 6A right).
-When neurons are equipped with homeostasis alone, the activation of sharp neuromodulation for the gastric mill rhythm leads to a rapid loss of function due to the unreliability of this combination (Fig 6B middle).
-However, with controlled neuromodulation, the gastric mill rhythm is successfully induced and sustained (Fig 6B right).
+Notably, variability was introduced in the initial conductance values across all neurons of the network to assess the robustness of the controllers in the presence of degeneracy (Fig 6B).
+When neurons are equipped with the calcium homeostatic controller only, the activation of sharp neuromodulation for the gastric mill rhythm leads to a rapid loss of function due to the unreliability of this combination (Fig 6C middle).
+However, with controlled neuromodulation, the gastric mill rhythm is successfully induced and sustained (Fig 6C right).
 Despite the intrinsic degeneracy of the neurons, reflected in significant variability in conductance ratios, the network achieves stable rhythmic output while maintaining appropriate intracellular calcium levels.
-This toy experiment demonstrates that the tandem of homeostasis and controlled neuromodulation can be applied in computational models to modulate and sustain not only single-neuron activity but also network-level dynamics.
+The mechanisms underlying these contrasting outcomes at the network level directly parallel those identified in the single-neuron analysis, as detailed below.
+More specifically, under sharp neuromodulation combined with calcium homeostasis (Fig 6C middle), the slow rhythm initially emerges but rapidly deteriorates as calcium-homeostatic compensation unfolds.
+As in the single-neuron case (Fig 4A), calcium homeostasis drives conductances along the native, unmodulated scaling direction, progressively pulling the slow-rhythm neurons (N4 and N5, green and yellow) away from the bursting isocline required to sustain anti-phase oscillations.
+The half-center oscillator mechanism critically depends on both neurons maintaining robust bursting with appropriate duty cycles.
+Consequently, the loss of bursting in even a single neuron is sufficient to collapse the entire slow rhythm.
+In contrast, under controlled neuromodulation (Fig 6C right), the activity-dependent feedback continuously corrects for homeostatic drift, maintaining each neuron on its target bursting isocline as described in Fig 4B.
+This ensures that the slow-rhythm neurons preserve the conductance ratios necessary for sustained anti-phase bursting, while the central neuron (N3, red) appropriately alternates between the fast and slow rhythmic patterns.
+Importantly, the initial conductance variability across all five neurons (Fig 6B) confirms that these network-level results are robust to degeneracy: despite substantial differences in individual conductance profiles, the tandem of controllers successfully coordinates population-wide rhythmic output.
+The fast rhythm (N1 and N2, purple and blue) remains active throughout, with only minor phase adjustments resulting from the electrical coupling to the slow-rhythm neurons.
+This is consistent with the experimental observation that pyloric rhythm modulation during gastric mill activation mainly involves changes in phase relationships [37], confirming that neuromodulation here primarily acts to activate or deactivate the slow rhythm.
+This toy experiment demonstrates that the tandem of calcium homeostasis and controlled neuromodulation can be applied in computational models to modulate and sustain not only single-neuron activity but also network-level dynamics.
 Discussion
-The efficient and robust integration of homeostasis and neuromodulation in neurons remains an open question, with several hypotheses proposed to explain their interaction [30, 6].
-Improper integration of these mechanisms – for instance, combining sharp neuromodulation with a homeostatic controller – can lead to negative interactions, resulting in unreliable behaviors or numerical instabilities.
-In contrast, pairing a homeostatic controller with controlled neuromodulation enables reliable modulation of neural function while maintaining appropriate calcium levels.
-This approach is biologically inspired, as neuromodulation is known to be activity-dependent, as observed experimentally [24, 45].
-Specifically, the neuromodulation controller models the role of G-protein-coupled receptors and second messengers, embedding feedback within the metabolic neuromodulation cascade [26, 46].
+The efficient and robust integration of calcium homeostasis and neuromodulation in neurons remains an open question, with several hypotheses proposed to explain their interaction [37, 7].
+Improper integration of these mechanisms, for instance combining sharp neuromodulation with a calcium-homeostatic controller, can lead to negative interactions, resulting in unreliable behaviors or numerical instabilities.
+In contrast, pairing a calcium-homeostatic controller with controlled neuromodulation enables reliable modulation of neural function while maintaining appropriate calcium levels.
+This approach is biologically inspired, as neuromodulation is known to be activity-dependent, as observed experimentally [30, 54, 19].
 This reliable interaction extends to neural circuits, such as the gastric mill circuit in crabs, where it contributes to robust network-level function [37].
@@
 As shown in Fig 4, it corresponds to the intersection of the strong bursting isocline (the target of controlled neuromodulation) and the homeostatic calcium setpoint.
-When no such intersection exists – e.g., under sodium channel blockade (Fig 5) – neither neuromodulated activity nor calcium homeostasis can be sustained.
+When no such intersection exists — e.g., under sodium channel blockade (Fig 5D) — neither neuromodulated activity nor calcium homeostasis can be sustained.
 This may lead to pathological compensations, where one controller prioritizes calcium regulation at the expense of the target firing pattern [59, 20, 2].
 Additionally, discontinuities in neuromodulated firing patterns may induce transient pathological behaviors.
+Moreover, the mere existence of such an intersection point does not guarantee that a neuron will reach it.
+For instance, to compensate for the loss of KCa channels (Fig 5C), a different neuromodulation target was required, corresponding to a different neuromodulator concentration.
+This means that, even when the intersection point exists, the neuron may require a specific neuromodulator concentration (or cocktail of neuromodulators) to partially or fully restore its function.
 Maximizing convergence between these control mechanisms requires increasing the probability of both an intersection and a continuous path in conductance space.
@@
 This highlights the crucial role of degeneracy in enabling robust, neuromodulated neural function.
-Finally, these findings underscore the potential risks associated with pharmacological ion channel blockade [5, 9, 21].
-If neuronal degeneracy is insufficient, such interventions may compromise the ability to sustain robust neuromodulation.
-As an alternative, we propose targeting elements within the neuromodulation cascade – such as second messengers or other components of the control loop – as more reliable pharmacological strategies.
+To further demonstrate the generality of the proposed framework, we applied it to a midbrain dopaminergic neuron model adapted from [52].
+As in the STG model, a transition from tonic spiking (pacemaking) to strong bursting was induced by modulating N-type and L-type calcium conductances ($\bar{g}_{\mathrm{CaN}}$ and $\bar{g}_{\mathrm{CaL}}$) through controlled neuromodulation.
+Calcium-homeostatic compensation continuously regulated calcium levels, ensuring a reliable and robust outcome despite initial conductance degeneracy within the neuronal population (see Supporting Information, S1 Appendix, for details).
+These constraints on compensation have direct implications for understanding neurological and neurodegenerative conditions.
+Our framework predicts that the clinical severity of channelopathies, in which mutations impair or eliminate specific ion channel functions, would depend on the degree of degeneracy of the affected channel [27]: mutations affecting highly degenerate channels may be largely compensated and produce mild phenotypes, whereas mutations affecting non-degenerate channels would lead to severe dysfunction.
+Similarly, the progressive loss of neuromodulatory input characteristic of neurodegenerative diseases, such as dopamine depletion in Parkinson’s disease [48], can be interpreted within our framework as a gradual loss of the controlled neuromodulation component.
+In this scenario, the calcium-homeostatic controller would increasingly operate alone, potentially producing the kind of pathological compensations observed in our sharp neuromodulation results (Fig 2A-C).
+This perspective suggests that therapeutic strategies aimed at preserving or restoring the activity-dependent feedback loop, rather than directly replacing the lost neuromodulator, may be more effective at maintaining robust neuronal function.
+These findings underscore the potential risks associated with pharmacological ion channel blockade [6, 10, 27].
+If neuronal degeneracy is insufficient (or if the amount of neuromodulator released is unchanged), such interventions may compromise the ability to sustain robust neuromodulation.
+As an alternative, we propose targeting elements within the neuromodulation cascade — such as second messengers or other components of the control loop — as more reliable pharmacological strategies.
 By preserving the dynamics of controlled neuromodulation, these interventions could offer a more robust alternative to direct ion channel blockade.
+Finally, the controlled neuromodulation framework presented here operates on single-compartment models.
+Extending it to multi-compartment models with dendritic conductances is a natural direction for future work.
+In principle, the DIC framework generalizes to spatially extended models, as the sensitivity matrix can be augmented to include dendritic conductances and electrotonic coupling terms [23].
+However, this extension raises challenges related to the increased dimensionality of the conductance space and the established non-uniformity of dendritic ion channel distributions [25].
+Furthermore, recent experimental evidence demonstrates that neuromodulators can act in a compartment-specific manner [67], and computational studies show that the same neuromodulatory action can have opposing effects at different dendritic locations [32].
+These observations suggest that spatially distributed or hierarchical control architectures may be required.
+Addressing these challenges in the context of controlled neuromodulation remains an open direction for future work.
 Materials and methods
@@
 Throughout this study, the isolated crab STG neuron model of [31] was employed.
-The STG model consists of seven ion channels that operate on various time scales: fast sodium channels ($\bar{g}_{\mathrm{Na}}$); delayed-rectifier potassium channels ($\bar{g}_{\mathrm{Kd}}$); T-type calcium channels ($\bar{g}_{\mathrm{CaT}}$); A-type potassium channels ($\bar{g}_{\mathrm{A}}$); slow calcium channels ($\bar{g}_{\mathrm{CaS}}$); calcium controlled potassium channels ($\bar{g}_{\mathrm{KCa}}$); and H channels ($\bar{g}_{\mathrm{H}}$).
-The homeostatic controller
-The homeostatic controller adjusts all conductances using an integration rule described in [38].
+The STG model consists of seven ion channels that operate on various time scales: fast sodium channels ($\bar{g}_{\mathrm{Na}}$); delayed-rectifier potassium channels ($\bar{g}_{\mathrm{Kd}}$); T-type calcium channels ($\bar{g}_{\mathrm{CaT}}$); A-type potassium channels ($\bar{g}_{\mathrm{A}}$); slow calcium channels ($\bar{g}_{\mathrm{CaS}}$); calcium-activated potassium channels ($\bar{g}_{\mathrm{KCa}}$); and H channels ($\bar{g}_{\mathrm{H}}$).
+The calcium-homeostatic controller
+The calcium-homeostatic controller adjusts all conductances using an integration rule described in [45].
 Let $\bar{g}_{i}$ denote the conductance of ion channel $i\in[1,N]$, and $m_{i}$ the corresponding mRNA concentration.
-The dynamics of the homeostatic controller are governed by the following equations:
+The dynamics of the calcium-homeostatic controller are governed by the following equations:
 $\tau_{i}\dot{m}_{i}$ $=[Ca^{+2}]_{\mathrm{target}}-[Ca^{+2}],$
@@
 Steady-state analysis reveals that the conductance ratios remain invariant, such that $\frac{\bar{g}_{i}}{\bar{g}_{j}}=\frac{\tau_{j}}{\tau_{i}}$.
-This invariance ensures that the relative scaling of conductances is preserved during homeostatic compensation.
+This invariance ensures that the relative scaling of conductances is preserved during calcium-homeostatic compensation.
 Consequently, the cell conductance properties develop along the homogeneous scaling beam defined by these ratios.
@@
 where $K_{p}$ and $K_{i}$ are the proportional and integral gains of the PI controller, respectively.
-In summary, the controller adjusts the modulated conductances $\bar{g}_{\mathrm{mod}}$ in response to the local concentration of neuromodulator, producing new reference values for these conductances to fine-tune the cell firing pattern – effectively neuromodulating it.
-This neuromodulator controller produces changes in conductance ratios that are normally preserved within the homeostatic controller.
-The new reference values depend on the neuromodulator concentration, the type of neuron, and the cell’s current state, characterized by the full set of conductances $\bar{g}_{\mathrm{ion}}\in\mathbb{R}^{N}$.
+In summary, the controller adjusts the modulated conductances $\bar{g}_{\mathrm{mod}}$ in response to the local concentration of neuromodulator, producing new reference values for these conductances to fine-tune the cell firing pattern — effectively neuromodulating it.
+This neuromodulator controller produces changes in conductance ratios that are normally preserved within the calcium-homeostatic controller.
+The new reference values depend on the neuromodulator concentration, the type of neuron, and the current state of the cell, characterized by the full set of conductances $\bar{g}_{\mathrm{ion}}\in\mathbb{R}^{N}$.
 The computation of these references is achieved through DICs, which encapsulate this complexity.